### v1

# Bayesian mixture modeling using a mixture of finite mixtures with normalized inverse Gaussian weights

Abstract
In Bayesian inference for mixture models with an unknown number of components, a finite mixture model is usually employed that assumes prior distributions for mixing weights and the number of components. This model is called a mixture of finite mixtures (MFM). As a prior distribution for the weights, a (symmetric) Dirichlet distribution is widely used for conjugacy and computational simplicity, while the selection of the concentration parameter influences the estimate of the number of components. In this paper, we focus on estimating the number of components. As a robust alternative Dirichlet weights, we present a method based on a mixture of finite mixtures with normalized inverse Gaussian weights. The motivation is similar to the use of normalized inverse Gaussian processes instead of Dirichlet processes for infinite mixture modeling. Introducing latent variables, the posterior computation is carried out using block Gibbs sampling without using the reversible jump algorithm. The performance of the proposed method is illustrated through some numerical experiments and real data examples, including clustering, density estimation, and community detection.
Keywords: Bayesian nonparametrics; Clustering; Dirichlet distribution; Inverse Gaussian distribution; Markov chain Monte Carlo; Mixture models
1 Introduction
We consider a finite mixture model with an unknown number of components:
$f(y|\theta_{1},\dots,\theta_{M},\pi)=\sum_{j=1}^{M}\pi_{j}f(y|\theta_{j}),$ (1.1)
where $M\in\{1,2,\dots\}$ is the number of components, $y\in\mathbb{R}^{d}$ is a $d$-dimensional observation vector, and $\pi=(\pi_{1},\dots,\pi_{M})^{\top}$ is the mixing weight vector, satisfying $\sum_{j=1}^{M}\pi_{j}=1$ and $\pi_{j}>0$. For $j=1,\dots,M$, $f(\cdot|\theta_{j})$ is a probability density or probability mass function parameterized by the component-specific parameter $\theta_{j}$. The function $f(\cdot|\theta_{j})$ is also called the kernel of the mixture model. Each observation is assumed to arise from one of the $M$ components, and each component is weighted based on its frequency of occurrence. Mixture models are important statistical models used in model-based clustering and density estimation because they can model complex data-generating distributions in random phenomena by using multiple components. Finite mixture models have been applied in many fields, for instance, sociology (Handcock et al., 2007), economics (Frühwirth-Schnatter and Kaufmann, 2008) and genetics (McLachlan et al., 2002) and so on. In application, the determination of $M$ is a very important issue. Correctly estimating $M$ improves model interpretability, the accuracy of kernel parameter estimates and model predictions, and reduces computation time. Although various methods have been proposed for the selection of $M$, such as model selection criteria and hypothesis testing, using these criterion makes it difficult to quantify the uncertainty of $M$, and also introduces the bias that model selection is carried out. The comprehensive review of (finite) mixture models is provided by McLachlan (2000); Frühwirth-Schnatter (2006); Frühwirth-Schnatter et al. (2019); McLachlan et al. (2019). In this paper, we focus on the case where $M$ is finite and make a clear distinction between the number of components $M$ and the number of clusters $k$. In general, the number of components is written as $M=k+M_{na}$, where $k$ is the number of components for which data are actually assigned, and $M_{na}$ is the number of empty components (see also Argiento and De Iorio, 2022).
In Bayesian analysis, a finite mixture model that assumes prior distributions for the mixing weights and the number of components is usually employed (see, e.g. Nobile, 1994; Miller and Harrison, 2018). Such a model is also called a mixture of finite mixtures (MFM), and the model is often used, as well as infinite mixture models represented by Dirichlet process mixture models. Although the reversible jump algorithm (Green, 1995; Richardson and Green, 1997) has been used to obtain samples from the posterior distribution based on MFM, it faces significant computational and implementation challenges. Methods based on marginal likelihoods have also been proposed (see, e.g. Nobile and Fearnside, 2007), while, as with reversible jump, the computational aspect is an issue. A method called sparse MFM (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016) has also been proposed, which focuses on estimating the number of clusters $k$ by using the overfitting model with a large fixed $M$, and choosing the prior distribution of the component rate well. Under same conditions, they showed that $k$ has consistency for true $M$. Since it allows for many empty components, we cannot estimate the number of components $M$ and it is not easy to quantify the uncertainty. Recently, the use of nonparametric Bayesian methods for MFM to estimate $M$ has attracted much attention. In Miller and Harrison (2018), the exchangeable partition distribution is derived by marginalizing out $M$, and the restaurant process is constructed to overcome computational difficulties. In Frühwirth-Schnatter et al. (2021), a generalized MFM in which the parameters of the Dirichlet distribution depend on $M$ is proposed. Miller and Harrison (2018) derived the theoretical result that the marginal posterior probabilities of $k$ and $M$ are asymptotically equivalent when the sample size is large, while only a sampling of $k$ is obtained as with general nonparametric Bayesian methods. Thus, the number of components $M$ cannot be sampled directly from the posterior distribution. Although Frühwirth-Schnatter et al. (2021)’s method can sample $M$ directly, as a sparse MFM, it has the disadvantage of producing a large number of empty components and overestimating $M$. In addition, most MFM studies use a symmetric Dirichlet distribution as the mixing weight vector, and the choice of a hyperparameter is very important. Although it is possible to set a prior distribution and learn from the data, the Metropolis-Hastings algorithm is required.
The main purposes of this paper are 1) sampling $M$ directly and efficiently compared to MFM based on the Dirichlet distribution; 2) estimating $M$ reasonably by suppressing $M_{na}$; 3) proposing a MFM that is robust to the choice of hyperparameters of the distribution on a simplex. To this end, we employ a normalized inverse Gaussian distribution as the mixing weight vector of the finite mixture model. To the best of our knowledge, there are few studies using the normalized inverse Gaussian distributions as the mixing weights in the field of finite mixture models. In Bayesian nonparametric inference with infinite mixtures, Lijoi et al. (2005) proposed a normalized inverse Gaussian process and showed some analytical results of the posterior distribution under the process. Similarly, since Miller and Harrison (2018) is a finite version of the Dirichlet process, the model presented in this paper corresponds to a kind of finite alternative of Lijoi et al. (2005). Furthermore, leveraging data augmentation with a latent gamma random variable and the result of Argiento and De Iorio (2022), we construct an efficient posterior sampling algorithm.
This paper is organized as follows. In Section 2, we introduce the MFM and its equivalent representation using discrete probability measures. We also discuss the data augmentation and the conditional posterior distribution of $M_{na}$, which are crucial for constructing the computational algorithm.In Section 3, we present the proposal method as well as the posterior computation algorithm. In Section 4, we illustrate some numerical studies to compare the proposed method with existing methods. R code implementing the proposed methods is available at Github repository:
https://github.com/Fumiya-Iwashige/MFM-Inv-Ga
2 Mixture of finite mixtures
We introduce a mixture of finite mixtures and its equivalent representation of the discrete probability measure. The proposed model and the posterior computation algorithm presented in the next section are largely based on the basic model presented in this section.
2.1 Formulation
Let each observations $Y_{1},\dots,Y_{n}$ be univariate or multivariate in an Euclidean space. A MFM model is a statistical model defined by the following hierarchical representation.
$\begin{split}&M-1\sim q_{M},\quad q_{M}\ \text{is a probability mass function (p.m.f.) on}\ \{0,1,2,\dots\},\\ &\pi|M\sim P_{\pi}(\pi|M),\\ &c_{i}|M,\pi\sim\mathrm{Categorical}_{M}(\pi),\\ &\xi_{m}|M\overset{\mathrm{i.i.d}}{\sim}p_{0}(\xi),\quad m=1,\dots,M,\\ &\theta_{i}|c_{i},\xi\overset{\mathrm{ind}}{\sim}\delta_{\xi_{c_{i}}}(\theta_{i}),\quad i=1,\dots n,\\ &Y_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}|\theta_{i}),\quad i=1,\dots,n,\end{split}$ (2.1)
where $f(y|\theta_{i})$ is a parametric model with parameter $\theta_{i}$, $\delta_{x}(\cdot)$ is a point mass at $x$ and $\xi$ is a element of a vector $\xi=(\xi_{1},\dots,\xi_{M})^{\top}$. $p_{0}(\xi)$ is a prior density function of the parameter in mixture components. The parameter space is denoted by $\Theta\subset\mathbb{R}^{d}$. Each $c_{i}$ is a latent allocation indicates to which component each observation is allocated. Given the number of components $M$, $P_{\pi}(\pi|M)$ is a probability distribution on the $M$-dimensional unit simplex, in that, this is the prior distribution for the mixing weights. We focus the case where the mixing weights can be further hierarchically expressed as follows:
$\begin{split}&S_{m}|M\overset{\mathrm{ind}}{\sim}h,\quad h\ \text{is a probability density function (p.d.f.) on}\ (0,\infty),\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi|M),\quad T=\sum_{m=1}^{M}S_{m},\end{split}$ (2.2)
where $S_{1},\dots,S_{M}$ is unnormalized weights and $\pi$ is (normalized) mixing weight vector. This representation is the most basic way to generate a probability distribution on the $d$-dimensional unit simplex $\mathbb{S}_{d}:=\{w=(w_{1},\dots,w_{d});w_{i}\geq 0,\sum_{i=1}^{d}w_{i}=1\}$. One of the most famous examples of a distribution on $\mathbb{S}_{d}$ is the Dirichlet distribution. If $S_{m}|M\sim\text{Gamma}(\gamma_{i},1)$ in (2.2), $\pi\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$, where $\text{Gamma}(a,b)$ is the gamma distribution with shape parameter $a>0$ and scale parameter $\beta>0$, and $\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$ is the Dirichlet distribution with parameter $\gamma=(\gamma_{1},\dots,\gamma_{d})>0$. If $\gamma_{1}=\cdots=\gamma_{d}=:\gamma$, the distribution obtained by normalization is $\text{Dirichlet}(\gamma,\dots,\gamma)$. The symmetric Dirichlet distribution is often used in the framework of MFM, because of conjugacy with categorical distributions and simplicity of computation. The symmetric structure of $\text{Dirichlet}(\gamma,\dots,\gamma)$ is also essential for marginalizing out $M$ and deriving the exchangeable partition distribution. However, it is known that the estimation result is sensitive to the choice of the shape parameter $\gamma$. For example, Miller and Harrison (2018) recommended to use $\gamma=1$ as a default choice and the value works well in many cases. The use of a small $\gamma$ was recommended in terms of sparse or generalized MFMs (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016; Frühwirth-Schnatter et al., 2021). A small value of $\gamma$ indicates that many empty components can be created. Although most previous studies allow for the appearance of the empty components, there are several problems. First, $M$ tends to be overestimated, making the interpretation of the model difficult. Second, the existing of the empty components decreases the predictive performance of the model, as we will see in a later section through density estimations. Finally, they lead to an increase in computation time. When we focus on the estimation of the number of components $M$, it is desirable to have few empty components. In other words, the discrepancy between $M$ and $k$ should be small. To achieve this, we use the normalized inverse Gaussian distribution as mixing weights, instead of the Dirichlet distribution.
2.2 Equivalent representations using discrete probability measures
We give an equivalent representation of the MFM. We can construct a discrete measure $P(\cdot)=\sum_{m=1}^{M}\pi_{m}\delta_{\xi_{m}}(\cdot)$ in the parameter space $\Theta$, almost surely, where $M$, $\xi$ and $\pi$ are realizations from the distributions $q_{M}$, $p_{0}$ and $P_{\pi}$, respectively. Let $\theta_{1},\dots,\theta_{n}$ be random variables according to $P$. Then, the model (2.1) is equivalent to the following hierarchical representation using a random measure $P$.
$\begin{split}&Y_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}|\theta_{i}),\quad i=1,\dots,n,\\ &\theta_{1},\dots,\theta_{n}|P\overset{\mathrm{i.i.d}}{\sim}P,\\ &P\sim\mathcal{P}(q_{M},h,p_{0}),\end{split}$ (2.3)
where $\mathcal{P}$ is a probability distribution of $P$ with parameters $q_{M}$, $h$ and $p_{0}$. The representation (2.3) is the MFM described as a nonparametric Bayesian framework with infinite mixtures. If we replace $\mathcal{P}$ with the Dirichlet process, (2.3) represents the well-known Dirichlet process mixture models (e.g., Escobar and West, 1995). If we replace $\mathcal{P}$ with the normalized inverse Gaussian process, (2.3) represents the normalized inverse Gaussian process mixture models (Lijoi et al., 2005). For MFM, Argiento and De Iorio (2022) proposed a normalized independent finite point process (Norm-IFPP), which is a class of flexible prior distributions for $P$. We employ the representation (2.3) with the Norm-IFPP. The advantages are as follows. We can directly estimate $M$ and $k$. This is a major difference from Miller and Harrison (2018), which estimates $M$ indirectly through the number of clusters $k$. Furthermore, an efficient Gibbs sampler can be constructed by incorporating the data augmentation with a latent Gamma random variable. This data augmentation enables us to overcome the lack of conjugacy in the categorical distribution. Instead of using the probability density function of $P_{\pi}$, we can use the density function of $h$. This is the key to building an efficient MCMC algorithm for the proposed model.
Our study is in the spirit of nonparametric Bayes in that it accounts for uncertainty in the prior distribution of the parameters by means of a random measure based on the Norm-IFPP. The details of Norm-IFPP and and the independent finite point process (IFPP) are given in Argiento and De Iorio (2022).
2.3 Data augmentation and conditional distribution of $M_{na}$
We illustrate the data augmentation and conditional posterior distribution of $M_{na}$. We introduce the data augmentation in models (2.1) and (2.2). This technique is employed in James et al. (2009) and Argiento and De Iorio (2022). Let $\mathcal{M}_{a}$ and $\mathcal{M}_{na}$ be the allocated and unallocated index sets, respectively. The conditional joint distribution of $\xi=(\xi_{1},\dots,\xi_{M})^{\top}$ and $S=(S_{1},\dots,S_{M})^{\top}$ given $M$ and a label vector $c$ is
$p(S,\tau|M,c)$ $\propto\left(\prod_{i=1}^{n}\frac{S_{c_{i}}}{T}\right)\left(\prod_{m=1}^{M}h(s_{m})p(\tau_{m})\right)$
$=\left(\prod_{m\in\mathcal{M}^{a}}\left(\frac{S_{m}}{T}\right)^{n_{m}}h(s_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}^{na}}h(s_{m})p(\tau_{m})\right),$
where $n_{m}=\#\{i:c_{i}=m\}$. Since this equation consists of categorical distributions, conjugacy with such a distribution is required as $P_{\pi}$ for Gibbs sampling. This restriction is relaxed by using latent a random variable $U_{n}|T\sim\mathrm{Gamma}(n,T)$, where $T=\sum_{m=1}^{M}S_{m}$. In fact,
$p(S,\tau,U_{n}|M,c)\propto u^{n-1}\left(\prod_{m\in\mathcal{M}}e^{-S_{m}u}S_{m}^{n_{m}}h(s_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}^{na}}e^{-S_{m}u}h(s_{m})p(\tau_{m})\right).$
Thus, the conditional distribution of each $S_{m}$ is
$S_{m}|U_{n},M,c\overset{\mathrm{ind}}{\sim}e^{-S_{m}u}S_{m}^{n_{m}}h(s_{m}),\quad m\in\mathcal{M}_{a},$ (2.4)
$S_{m}|U_{n},M,c\overset{\mathrm{i.i.d}}{\sim}e^{-S_{m}u}h(s_{m}),\quad m\in\mathcal{M}_{na}.$ (2.5)
If it is easy to generate random variables from the distributions in (2.4) and (2.5), an efficient Gibbs sampling algorithm can be constructed. For example, when the prior distribution of the unnormalized weight is an inverse Gaussian distribution, (2.4) and (2.5) are generalized inverse Gaussian distributions. Introducing $U_{n}$, the update of the variable $\pi$ is replaced by the update of the variable $S_{m}$. Thus, the selection of $h$ is essential in MCMC updates.
Under this data augmentation, the conditional distribution of $M_{na}$ is established in Theorem 5.1 in Argiento and De Iorio (2022). This theorem states that if $\mathcal{P}(q_{M},h,p_{0})$ follows a Norm-IFPP, then the posterior distribution of the random measure $P$, given $U_{n}$, is a superposition of a finite point process with fixed points and an IFPP. The IFPP characterizes the process of unallocated jumps, where the discrete probability distribution that serves as its parameter corresponds to the distribution of $M_{na}$. The conditional distribution of $M_{na}$ is given by
$\mathbb{P}(M_{na}=m|\theta,U_{n}=u)\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=1,2,\dots,$ (2.6)
where $\theta=(\theta_{1},\dots,\theta_{n})^{\top}$, $k$ is the number of cluster (unique values of $\theta$) and $\psi$ is a Laplace transform of $h$. We sample $M_{na}$ from (2.6) in the MCMC algorithm. Then, we straightforwardly obtain $M$ by adding $k$ to $M_{na}$.
Remark 2.1.
Although we use the data augmentation $U_{n}|T\sim\text{Gamma}(n,T)$, the distributions (2.4) and (2.5) are generally complex. In addition, the result of Argiento and De Iorio (2022) applies only to $h$ that have a Laplace transform. Thus, the inverse Gaussian distribution is one of the few examples that satisfy both computational and theoretical constraints.
3 Methodology
In this section, we propose a mixture of finite mixtures with normalized inverse Gaussian weights. Moreover, we develop an efficient posterior sampling algorithm based on Argiento and De Iorio (2022).
3.1 Mixture of finite mixtures with normalized inverse Gaussian weights
We propose a mixture of finite mixtures with normalized inverse Gaussian weights (denoted by MFM-Inv-Ga), where the notation explicitly specifies $h$ because it is essential for computation. Our proposal model only requires (2.2) to be
$S_{m}|M\overset{\mathrm{ind}}{\sim}\text{Inv-Ga}(\alpha,1),\quad\alpha>0,$
$\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim\text{Norm-Inv-Ga}(\alpha,\dots,\alpha),$
[…]
$h(s)=\frac{\alpha}{\sqrt{2\pi}}s^{-3/2}\exp\left(-\frac{1}{2}\left(\frac{\alpha^{2}}{s}+\beta^{2}s\right)+\beta\alpha\right),$ (3.1)
where $\alpha>0$ is the shape parameter and $\beta>0$ is the scale parameter. The mean and variance of $S$ are given by
$\mathbb{E}[S]=\frac{\alpha}{\beta},\quad\mathrm{Var}\left[S\right]=\frac{\alpha}{\beta^{3}}.$ (3.2)
From Lijoi et al. (2005), the probability density function of the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0,\ i=1,\dots,d$ is given by
$\begin{split}f(\pi)=&\frac{e^{\sum_{i=1}^{d}\alpha_{i}}\prod_{i=1}^{d}\alpha_{i}}{2^{d/2-1}\pi^{d/2}}\times K_{-d/2}\left(\sqrt{\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})}\right)\\ &\times\left(\pi_{1}^{3/2}\cdots\pi_{d-1}^{3/2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{3/2}\times\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})^{d/4}\right)^{-1},\end{split}$ (3.3)
where $\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})=\sum_{i=1}^{d-1}\alpha_{i}^{2}/\pi_{i}+\alpha_{d}^{2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{-1}$ and $K_{m}$ is the modified Bessel functions of the third kind of order $m$ (see, also Ghosal and van der Vaart, 2017). Due to the breakdown of conjugacy with the categorical distribution, constructing an efficient Gibbs sampler is challenging. Because of the complexity of (3.3), the calculation in Miller and Harrison (2018) is intractable: constructing an MCMC algorithm based on the restaurant process by marginalizing out $M$ is extremely challenging. To overcome this difficulty, we use the data augmentation in Subsection 2.3 and the method proposed in Argiento and De Iorio (2022). As a result, it is sufficient to use not (3.3) but (3.1) and (3.8) to construct our MCMC algorithm.
3.2 Posterior computation
We present a fast and efficient posterior computation algorithm for the proposed method. To this end, we adopt the blocked Gibbs sampling scheme in Argiento and De Iorio (2022). Let $\eta$ be a hyper-parameter of the prior distribution $q_{M}$, and $p(\cdot)$ be a joint prior density function except for $h$ and $q_{M}$. $n_{m}=\#\{i;c_{i}=m\}$ is the size of the $m$th cluster. $\psi(u)$ is the Laplace transform of $h$, which is defined by $\psi(u):=\mathbb{E}[e^{-uS_{m}}]$ for $u\geq 0$.
Algorithm 1 Block Gibbs sampler
  Step 0. Set initial values.
  Step 1. Sample $U_{n}$ from $\mathrm{Gamma}(n,T)$, where $T=\sum_{m=1}^{M}S_{m}$.
  Step 2. Sample $c_{i},for\ i=1,\dots,n,$ the following discrete probability distribution defined by each probabilities
$\mathbb{P}(c_{i}=j|\mathrm{rest})\propto S_{j}f(y_{i}|\tau_{j}),\quad j=1,\dots,M.$
  Step 3. Sample the hyper parameter $\eta$ of the $q_{M}$ from
$p(\eta|\mathrm{rest})\propto\Psi(u,k)p(\eta),\quad\Psi(u,k):=\sum_{m=0}^{\infty}\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k).$
  Step 4. Sample $M_{na}$ from the following discrete distribution
$\mathbb{P}(M_{na}=m|\mathrm{rest})\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=0,1,\dots.$
[…]
$p(\tau_{m}|\mathrm{rest})\propto\left\{\prod_{i;\theta_{i}=\tau_{m}}f(y_{i}|\tau_{m})\right\}p(\tau).$
  Step 7. Sample $S_{m_{na}}$, for $\ m_{na}=k+1,\dots,M_{na}+k$from $\mathrm{GIG}(2u+1,\alpha^{2},-1/2).$
[…]
3.3 Prior distributions for the number of mixture components
Assume that $M-1$ follows a discrete probability distribution with the support $\{0,1,2,\dots\}$. In this paper, we consider $\text{Poisson}(\Lambda)$ ($\Lambda>0$) for $M-1$, because the constraints in Steps 3 and 4 are satisfied. In fact, from Argiento and De Iorio (2022), we have
$\Psi(u,k)=\Lambda^{k-1}(\Lambda\psi(u)+k)\exp(\Lambda(\psi(u)-1)).$ (3.4)
Assuming $\Lambda\sim\text{Gamma}(a_{\Lambda},b_{\Lambda})$ for $a_{\Lambda},b_{\Lambda}>0$, the full conditional distributions of $\Lambda$ and $M_{na}$ are given by
$p(\Lambda|\mathrm{rest})\propto\psi(u)(k+a_{\Lambda}-1)$ $\mathrm{Ga}(k+a_{\Lambda}+1,1-\psi(u)+b_{\Lambda})$
$+k(b_{\Lambda}+1-\psi(u))\mathrm{Ga}(k+a_{\Lambda},1-\psi(u)+b_{\Lambda}),$
[…]
3.4 Specification of kernels
The choice of kernel is important, and the appropriate kernel must be selected for the purpose. In this paper, although we do not discuss the details of the selection of kernels, we present some famous kernels for the sake of completeness.
3.4.1 Cluster analysis and density estimation
One of the most famous and useful kernels is the (multivariate) normal kernel $N(\cdot|\mu,\sigma^{2})$. In the univariate case, we often use the normal inverse gamma model $N(\mu|m_{0},\sigma^{2}/\eta)\times\text{IG}(\sigma^{2}|c_{0},C_{0})$ as a prior distribution of $\xi=(\mu,\sigma^{2})$, where $\ m_{0}\in\mathbb{R}$, $\eta,\ c_{0},\ C_{0}>0$. The parameter $\eta$ is called a smoothing parameter and plays an important role in density estimation. It is possible to include a hierarchical prior for $\eta$. In the multivariate case, we often employ the normal inverse Wishart model $N_{r}(\mu|m_{0},\Sigma/B_{0})\times\mathrm{W}^{-1}(\Sigma|c_{0},C_{0})$ as a prior distribution of $(\mu,\Sigma)$, where $\ m_{0}\in\mathbb{R}^{r},\ B_{0}>0,\ c_{0}>r-1$ and $C_{0}$ is a positive definite matrix. We will use normal kernels in later numerical experiments in Sections 4.1 and 4.2 for clustering and density estimation. In the context of finite mixture models, various kernels have been proposed. If we have prior information that the data have skewness, the skew normal or skew-t kernel is also useful (see e.g., Frühwirth-Schnatter and Pyne, 2010). The skew-normal and skew-t distributions is easy to handle because these kernels have the scale mixtures of normal representations, except for the degree of freedom of skew-t distribution.
3.4.2 Network analysis
As an application of the proposed method, we perform community detection on network data. Community detection is the task of identifying dense subclasses in network data and corresponds to clustering and estimating the number of components, called the number of communities in network analysis. Note that the number of components of finite mixture models is equivalent to the number of communities in the network, and both are denoted $M$. Estimating the number of communities is an important problem and various methods have been proposed (Shi and Malik, 2000; Newman, 2004; White and Smyth, 2005). From a model-based perspective, it is essentially the same as estimating the number of components in a finite-mixture model, and we can apply MFM. The stochastic block model is a famous statistical model of network data (Henze, 1986; Nowicki and Snijders, 2001; Geng et al., 2019), which assumes a stochastic block structure behind and specifies the community structure by estimating the probability of edges being drawn between each group. We estimate the number of communities with the MFM in the framework of this stochastic block model. Geng et al. (2019) proposed a stochastic block model based on MFM with Dirichlet weights, and also constructed a similar algorithm to Miller and Harrison (2018).
MFM can be easily applied to community detection by modifying the kernel. Data $y$ are replaced by the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $A=(A_{ij})\in\{0,1\}^{n\times n}$ and $n$ is the size of the node. When $A_{ij}=1$, this indicates that an edge is drawn from the $i$ th node to the $j$ th node, and when $A_{ij}=0$, this does not. For simplicity, we assume that the adjacency matrix is not direct and does not have a self-loop, in that $A_{i,j}=A_{ji}$ and $A_{ii}=0$, where $1\leq i<j\leq n$. The stochastic block model is formulated as follows,
$\begin{split}&A_{ij}|Q,M\sim\text{Bernoulli}(\theta_{ij}),\quad\theta_{ij}=Q_{c_{i}c_{j}},\quad 1\leq i<j\leq n,\\ &Q_{rs}\sim\text{Beta}(a_{Q},b_{Q}),\quad 1\leq r\leq s\leq M,\end{split}$ (3.5)
where $a_{Q},\ b_{Q}>0$ and $M$ is the number of communities. $Q=(Q_{rs})\in[0,1]^{M\times M}$ is a symmetric matrix and defines the stochastic block structure of a network. Each element $Q_{rs}$ represents the probability that an edge is drawn between any node belonging to the community label $r$ and any node belonging to the community label $s$. To perform community detection based on the proposed model, we just set the Bernoulli likelihood as the kernel.
3.5 Evaluation of the number of empty components
From the point of view of the interpretability of the model, the generalization performance, and the computational efficiency, it is reasonable that the number $M_{na}$ should be small. For the full conditional distribution of $M_{na}$, the following inequality holds, where $M-1|\Lambda\sim\mathrm{Poisson}(\Lambda)\ \text{and}\ \Lambda\sim\mathrm{Gamma}(a_{\Lambda},b_{\Lambda})\ \text{for}\ a_{\Lambda},b_{\Lambda}>0$.
Proposition 3.1.
For the full conditional distribution of $M_{na}$, the inequalities
$\mathbb{P}(M_{na}\geq 1|U_{n}=u,\ k,\ M=m,\ \Lambda)$ $\leq\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right),$ (3.6)
$\mathbb{P}(M_{na}\geq 1|U_{n}=u,\ k,\ M=m)$ $\leq\psi(u)\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right)$ (3.7)
holds, where $\psi$ is the Laplace transform of the probability distribution for $h$.
Proof.
When $m\neq 0$,
$\mathbb{P}(M_{na}=m|U_{n}=u,\ k,\ M=m,\ \Lambda)=\frac{k}{\Lambda\psi(u)+k}\mathcal{P}_{0}(m;\Lambda\psi(u))+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\mathcal{P}_{1}(m;\Lambda\psi(u)).$
Thus, the conditional expectation of $M_{na}$ is
$\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m,\ \Lambda]$ $=\frac{k}{\Lambda\psi(u)+k}\times\Lambda\psi(u)+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\times(\Lambda\psi(u)+1)$
$=\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right).$
Furthermore,
$\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m]$ $=\mathbb{E}_{\Lambda}[\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m,\ \Lambda]]$
$=\int\left(\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right)\right)\frac{b_{\Lambda}^{a_{\Lambda}}}{\Gamma(a_{\Lambda})}\Lambda^{a_{\Lambda}-1}\exp(-b_{\Lambda}\Lambda)\mathrm{d}\Lambda$
$\leq\psi(u)\times\frac{a_{\Lambda}}{b_{\Lambda}}+\frac{\psi(u)}{k}\times\frac{a_{\Lambda}}{b_{\Lambda}}$
$=\psi(u)\times\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right).$
Finally, the result follows from Markov’s inequality. ∎
[…]
$\psi_{\text{Inv-Ga}}(u)=\exp\left(\alpha\left(1-\sqrt{1+2u}\right)\right),\quad u\geq 0,$ (3.8)
$\psi_{\text{Ga}}(u)=\left(\frac{\eta}{\eta+u}\right)^{\gamma},\quad u\geq 0.$ (3.9)
Laplace transforms $\psi_{\text{Inv-Ga}}$ and $\psi_{\text{Ga}}$ are decreasing functions with respect to $u$. The former has exponential decay, while the latter is polynomial. Figure 1 shows the graphs of the Laplace transforms of inverse Gaussian and gamma distributions when the shape parameters are $\alpha,\gamma=1.0,0.2,10^{-1},10^{-2},10^{-3}$ and the scale parameters are $1$. It can be seen that $\psi_{\text{Inv-Ga}}$ decreases much faster than $\psi_{\text{Ga}}$. The speed of this decrease has a significant impact on the appearance of the empty component in estimating the number of components.
[…]
Inequalities (3.6) and (3.7) are not for marginal posterior distributions, but for full conditional distributions of $M_{na}$. This is due to the fact that the marginal posterior distributions of $U_{n}$ are analytically intractable.
4 Empirical demonstrations
We evaluate the performance of the MFM-Inv-Ga and MFM-Ga methods through some numerical experiments. Recall that $\gamma$ is the shape parameter of the prior distribution of unnormalized weights in the proposed method (MFM-Inv-Ga), while $\alpha$ is that of MFM-Ga.
4.1 Inference for the number of mixture components and clustering
In this subsection, we illustrate the performance of clustering and inference for the number of components using artificial and real data.
4.1.1 Artificial data
In this simulation, we assume that $M_{\mathrm{true}}=3$. We generate data from the following multivariate normal distribution:
$f(y|\mu_{1},\mu_{2},\mu_{3})=0.8\cdot N_{2}(y|\mu_{1},I_{2})+0.1\cdot N_{2}(y|\mu_{2},I_{2})+0.1\cdot N_{2}(y|\mu_{3},I_{2}),$
[…]
  • The posterior mean of the number of components $M$.
  • The posterior mean of the rand index. The rand index is a measure of the clustering fitting, and the value takes $[0,1]$. When it is close to $1$, the assignment estimate is reasonable.
  • The posterior probability that the number of empty components $M_{na}$ is equal to 00.
The respective averages over $50$ repetitions are denoted by $\widehat{\mathrm{RI}}$ and ^$\widehat{\mathbb{P}}(M_{na}=0)$.
We report the posterior mean of $M$ in Figure 2. It is observed that the results of the MFM-Inv-Ga method remain almost the same even if the shape parameter is varied. However, the MFM-Ga method tends to overestimate the number of mixture components as the shape parameter decreases.
Figure 2: Box plots of the posterior mean of the number of components $M$ for each shape parameter. From left to right, shape parameters are set as $\gamma,\alpha=1,0.2,10^{-1},10^{-2},10^{-3}$.
Table 1 shows the results of the clustering performance and the posterior probability that $M_{na}$ is equal to 00. For clustering, the MFM-Inv-Ga and MFM-Ga methods are comparable. Both methods have reasonable clustering accuracy, only slightly better for MFM-Ga than for MFM-Inv-Ga. As the shape parameters $\alpha$ and $\gamma$ decrease, $\widehat{\mathrm{RI}}_{i}$ increases slightly and its standard deviation decreases. This is natural given that for data that have components with relatively very large cluster sizes, a suitable prior on a simplex is one with large mass at the edges or vertices. It is important to note that Figure 2 and Table 1 show that the MFM-Inv-Ga method produces few empty components for all scenarios and provides reasonable estimates of $M$, but not MFM-Ga.
Table 1: Results of the rand index $\widehat{\mathrm{RI}}_{i}$ and posterior probability that the number of empty components is equal to zero ^$\widehat{\mathbb{P}}(M_{na}=0)$ (their standard deviations are shown in parentheses) averaged over 50 Monte Carlo replications for each shape parameters.
$\widehat{\mathrm{RI}}_{i}$
$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
MFM-Inv-Ga $0.993$ $0.994$ $0.995$ $0.995$ $0.995$
$(0.013)$ $(0.006)$ $(0.006)$ $(0.004)$ $(0.005)$
MFM-Ga $0.995$ $0.995$ $0.995$ $0.996$ $0.996$
$(0.006)$ $(0.004)$ $(0.003)$ $(0.004)$ $(0.003)$
^$\widehat{\mathbb{P}}(M_{na}=0)$
$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
MFM-Inv-Ga $1.000$ $0.995$ $0.982$ $0.947$ $0.952$
$(0.001)$ $(0.008)$ $(0.032)$ $(0.135)$ $(0.087)$
MFM-Ga $0.979$ $0.555$ $0.328$ $0.088$ $0.066$
$(0.008)$ $(0.080)$ $(0.066)$ $(0.012)$ $(0.013)$
This can be seen in Table 2. It can also be seen that MFM-Inv-Ga assigns a very high posterior probability to $M_{\mathrm{true}}$ than MFM-Ga, and the behavior of the posterior distributions for $k$ and $M$ is similar, together with Table 1.
Table 2: Posterior probabilities and their standard deviations (shown in parentheses) of the number of components $M$ averaged over 50 Monte Carlo replications. The highest value is bolded.
MFM-Inv-Ga
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
$\alpha=1$ $0.000$ $0.000$ $\bf{0.973}$ $0.027$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.093)$ $(0.093)$ $(0.000)$ $(0.000)$
$\alpha=0.2$ $0.000$ $0.080$ $\bf{0.886}$ $0.033$ $0.001$ $0.000$
$(0.000)$ $(0.274)$ $(0.268)$ $(0.048)$ $(0.003)$ $(0.000)$
$\alpha=10^{-1}$ $0.000$ $0.000$ $\bf{0.942}$ $0.051$ $0.006$ $0.001$
$(0.000)$ $(0.000)$ $(0.084)$ $(0.069)$ $(0.013)$ $(0.005)$
$\alpha=10^{-2}$ $0.000$ $0.020$ $\bf{0.890}$ $0.063$ $0.015$ $0.012$
$(0.000)$ $(0.141)$ $(0.199)$ $(0.078)$ $(0.039)$ $(0.048)$
$\alpha=10^{-3}$ $0.000$ $0.050$ $\bf{0.860}$ $0.068$ $0.015$ $0.007$
$(0.000)$ $(0.209)$ $(0.225)$ $(0.084)$ $(0.029)$ $(0.020)$
MFM-Ga
$\gamma=1$ $0.000$ $0.000$ $\bf{0.948}$ $0.049$ $0.002$ $0.000$
$(0.000)$ $(0.000)$ $(0.055)$ $(0.049)$ $(0.006)$ $(0.001)$
$\gamma=0.2$ $0.000$ $0.010$ $\bf{0.508}$ $0.301$ $0.121$ $0.061$
$(0.000)$ $(0.070)$ $(0.093)$ $(0.037)$ $(0.037)$ $(0.035)$
$\gamma=10^{-1}$ $0.000$ $0.000$ $\bf{0.307}$ $0.306$ $0.196$ $0.191$
$(0.000)$ $(0.007)$ $(0.093)$ $(0.037)$ $(0.037)$ $(0.035)$
$\gamma=10^{-2}$ $0.000$ $0.003$ $0.088$ $0.155$ $\bf{0.176}$ $0.578$
$(0.000)$ $(0.002)$ $(0.021)$ $(0.016)$ $(0.013)$ $(0.071)$
$\gamma=10^{-3}$ $0.000$ $0.003$ $0.066$ $0.132$ $\bf{0.163}$ $0.636$
$(0.000)$ $(0.020)$ $(0.019)$ $(0.013)$ $(0.011)$ $(0.072)$
Table 3 shows the CPU times of MFM-Inv-Ga and MFM-Ga for $50$ repetitions of MCMC with $2000$ iterations, where $\alpha,\gamma=1.0$ and $10^{-3}$. When $\alpha=\gamma=1.0$, MFM-Inv-Ga is faster than MFM-Ga on average, but MFM-Inv-Ga has more variability than MFM-Ga. On the other hand, MFM-Ga is very time-consuming in $\gamma=10^{-3}$, because many empty components are created. As a result, MFM-Inv-Ga has the same clustering accuracy as MFM-Ga and outperforms MFM-Ga in terms of $M$ estimation and CPU time.
Table 3: Comparison of CPU times (in seconds) and their standard deviations (shown in parentheses) for artificial data averaged over $50$ replications, $\alpha,\ \gamma=1.0,10^{-3}$
MFM-Inv-Ga MFM-Ga
$\alpha=\gamma=1$ $88.545$ $90.898$
$(5.686)$ $(2.774)$
$\alpha=\gamma=10^{-3}$ $87.061$ $192.234$
$(12.286)$ $(8.932)$
4.1.2 Thyroid Data
We apply the proposed method to famous thyroid data. The data is available from the R package mclust and is well known as benchmark data for clustering. The sample size is $215$ and the dimension of the data is $6$. The thyroid disease of each patient is included and classified into three categories: Normal, hypo and hyper. The number of diseases is $150$, $30$ and $35$, respectively. Using these labels as true labels, the main interest is the accuracy of the clustering and whether the number of components is estimated to be $3$.
We use the same model and shape parameters as in Section 4.1.1. We independently run the $5$ MCMC chain with different initial values. The number of iterations for each chain is $20000$ and finally we get the $10000$ samples. We compare MFM-Inv-Ga and MFM-Ga through the posterior mean of $M$, the posterior probability $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ and the posterior mean of the rand index (RI). Table 4 shows that the result of the real data analysis is similar to that of Section 4.1.1. MFM-Inv-Ga provides a reasonable estimate of $M$ for all shape parameters by not producing empty components. However, MFM-Ga overestimates $M$. Furthermore, MFM-Inv-Ga is slightly more accurate than MFM-Ga.
Table 4: Posterior means of $M$ ($\hat{M}$), posterior probabilities of $M_{na}=0$ ($\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$) and posterior means of rand index (RI) for thyroid data.
MFM-Inv-Ga
$\hat{M}$ $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ RI
$\alpha=1$ $3.200$ $1.000$ $0.895$
$\alpha=0.2$ $3.213$ $0.994$ $0.895$
$\alpha=10^{-1}$ $3.225$ $0.987$ $0.894$
$\alpha=10^{-2}$ $3.222$ $0.987$ $0.895$
$\alpha=10^{-3}$ $3.489$ $0.960$ $0.893$
MFM-Ga
$\hat{M}$ $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ RI
$\gamma=1$ $3.670$ $0.962$ $0.893$
$\gamma=0.2$ $4.944$ $0.434$ $0.890$
$\gamma=10^{-1}$ $4.945$ $0.394$ $0.893$
$\gamma=10^{-2}$ $7.425$ $0.067$ $0.893$
$\gamma=10^{-3}$ $7.626$ $0.046$ $0.895$
4.2 Density estimation
As seen in the previous subsection, the main difference between the MFM-Inv-Ga and MFM-Ga methods is the frequency of occurrence of the empty components. The MFM-Ga method can achieve very high clustering accuracy by setting a small $\gamma$ and allowing empty components. However, small $\gamma$ not only makes the model difficult to interpret but also degrades the predictive accuracy of the model. In this subsection, we examine this phenomenon through density estimation using predictive distributions.
We used the famous galaxy dataset, which is a small data set consisting of $82$ velocities (km/sec) of different galaxies. The data is widely used in nonparametric Bayesian statistics as a benchmark for density estimation and cluster analysis. The details of the data are found in Roeder (1990). We set $M-1|\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1/5)$. We also employ the univariate normal kernel $N(y|\mu,\sigma^{2})$, and the normal–inverse gamma conjugate prior $N(\mu|m_{0},\sigma^{2}/\tau)\times\text{IG}(\sigma^{2}|c_{0},C_{0})$ as the prior of the parameters in the kernel. Moreover, we assume that the prior of $C_{0}$ is $\text{Gamma}(d_{0},D_{0})$, and we set $m_{0}=(\max(\mathrm{data})+\min(\mathrm{data}))/2,\ d_{0}=0.2,\ D_{0}=10/(\max(\mathrm{data})+\min(\mathrm{data}))^{2}$ as Richardson and Green (1997). The parameter $\tau$ controls the smoothness of the estimated density function. We assume that the prior of the smoothing parameter $\tau$ is $\text{Gamma}(w,W)$, where $w=0.5$ and $W=50$ as Escobar and West (1995). Parameters $\tau$ and $C_{0}$ should be carefully learned from the data, since density estimation is sensitive to their choice. The MCMC iterations are $100000$ and the first half of the $90000$ samples are not used as a burn-in period. The shape parameters are the same as in Section 4.1. We evaluated the result of density estimation and posterior probabilities of $k$, $M$, and $M_{na}$ for each shape parameter.
Figure 3: Results of density estimation for galaxy data using MFM-Inv-Ga (left) and MFM-Ga (right).
We show the estimated density functions using the posterior means in Figure 3. From the figure, it is observed that the shapes of the estimated densities using the MFM-Inv-Ga do not depend on the choice of the shape parameter $\alpha$. However, results using the MFM-Ga method seem to be strongly influenced by the choice of the shape parameter $\gamma$. For $\gamma=10^{-2}$ and $10^{-3}$, MFM-Ga cannot capture two large peaks in the middle.
Table 5: Posterior probabilities of the number of components $M$ and clusters $k$ for Galaxy data.
MFM-Inv-Ga
$M\leq 3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M=9$ $M\geq 10$
$\alpha=1.0$ $0.000$ $0.089$ $0.224$ $0.31$ $0.170$ $0.097$ $0.054$ $0.056$
$\alpha=0.2$ $0.136$ $0.194$ $0.171$ $0.152$ $0.114$ $0.086$ $0.054$ $0.091$
$\alpha=10^{-1}$ $0.071$ $0.096$ $0.154$ $0.144$ $0.137$ $0.102$ $0.079$ $0.216$
$\alpha=10^{-2}$ $0.039$ $0.078$ $0.119$ $0.151$ $0.123$ $0.102$ $0.081$ $0.308$
$\alpha=10^{-3}$ $0.057$ $0.111$ $0.170$ $0.153$ $0.122$ $0.101$ $0.071$ $0.217$
$k\leq 3$ $k=4$ $k=5$ $k=6$ $k=7$ $k=8$ $k=9$ $k\geq 10$
$\alpha=1.0$ $0.000$ $0.092$ $0.241$ $0.331$ $0.164$ $0.092$ $0.046$ $0.034$
$\alpha=0.2$ $0.156$ $0.215$ $0.206$ $0.182$ $0.118$ $0.063$ $0.033$ $0.026$
$\alpha=10^{-1}$ $0.089$ $0.127$ $0.214$ $0.224$ $0.169$ $0.094$ $0.052$ $0.031$
$\alpha=10^{-2}$ $0.088$ $0.123$ $0.201$ $0.232$ $0.160$ $0.107$ $0.054$ $0.035$
$\alpha=10^{-3}$ $0.082$ $0.152$ $0.256$ $0.221$ $0.143$ $0.081$ $0.039$ $0.025$
MFM-Ga
$M\leq 3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M=9$ $M\geq 10$
$\gamma=1.0$ $0.038$ $0.104$ $0.191$ $0.206$ $0.176$ $0.122$ $0.072$ $0.091$
$\gamma=0.2$ $0.018$ $0.042$ $0.076$ $0.098$ $0.114$ $0.117$ $0.103$ $0.432$
$\gamma=10^{-1}$ $0.019$ $0.040$ $0.057$ $0.071$ $0.075$ $0.083$ $0.079$ $0.577$
$\gamma=10^{-2}$ $0.009$ $0.023$ $0.042$ $0.059$ $0.071$ $0.076$ $0.075$ $0.645$
$\gamma=10^{-3}$ $0.003$ $0.004$ $0.013$ $0.019$ $0.025$ $0.033$ $0.038$ $0.865$
$k\leq 3$ $k=4$ $k=5$ $k=6$ $k=7$ $k=8$ $k=9$ $k\geq 10$
$\gamma=1.0$ $0.044$ $0.120$ $0.232$ $0.236$ $0.188$ $0.104$ $0.044$ $0.033$
$\gamma=0.2$ $0.145$ $0.198$ $0.304$ $0.186$ $0.098$ $0.040$ $0.019$ $0.009$
$\gamma=10^{-1}$ $0.145$ $0.198$ $0.304$ $0.186$ $0.098$ $0.040$ $0.019$ $0.009$
$\gamma=10^{-2}$ $0.775$ $0.195$ $0.024$ $0.005$ $0.000$ $0.000$ $0.000$ $0.000$
$\gamma=10^{-3}$ $0.983$ $0.018$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
The number of clusters in the galaxy data has been reported as $5$ or $6$ in existing studies. From table 5, in MFM-Inv-Ga, the posterior distributions of $k$ have large probabilities at $k=5$ and $6$ for all $\alpha$ and each posterior distribution induced by MFM-Inv-Ga is more similar than by MFM-Ga. For $M$, both MFM-Inv-Ga and MFM-Ga overestimate when $\alpha$ and $\gamma$ are small. This can be seen in Table 6. The reason is why the data size $n=82$ is small and $M_{na}$ is easily produced. Tables 5 and 6 show that the degree of overestimation is much smaller for MFM-Inv-Ga than for MFM-Ga.
Table 6: Posterior probabilities of the number of empty components is equal to zero for Galaxy data
$\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$
$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
MFM-Inv-Ga $0.854$ $0.654$ $0.480$ $0.347$ $0.443$
MFM-Ga $0.648$ $0.081$ $0.040$ $0.010$ $0.002$
Focusing on the number of clusters $k$, it is interesting that MFM-Inv-Ga is more robust for shape parameter estimates $k$ than MFM-Ga. This suggests that the prior distribution of $k$ based on MFM-Inv-Ga is less informative than MFM-Ga, i.e., the same relationship holds for MFM-Inv-Ga and MFM-Ga as for normalized inverse Gaussian process and Dirichlet process. In Lijoi et al. (2005), the prior of $k$ induced by a normalized inverse Gaussian process is not more informative for the precision parameter than the Dirichlet process. However, the prior $k$ for the normalized inverse Gaussian process can be written in closed form, while that for MFM-Inv-Ga is given in complex integral form.
The shape parameter of MFM-Ga should be chosen carefully because it has a significant impact on clustering, density estimation, and the appearance of empty components. However, MFM-Inv-Ga is much more robust than MFM-Ga with respect to the choice of the shape parameter. Hence, MFM-Inv-Ga is superior to MFM-Ga in that it is much easier to use than MFM-Ga.
4.3 Community detection
We apply the proposed method to the community detection for network data. Similar comparisons as in Section 4.1 are made for both artificial and real data. Since the number of components of the finite mixture models is equivalent to the number of communities of the network, we use the same notation $M$ to denote the number of communities.
4.3.1 Artificial data
First, we illustrate the performance of the proposed method using simulation data. We assume that the true number of communities is $3$, denoted by $M_{\mathrm{true}}=3$ and the number of nodes in the network is set as $n=150$. We here consider the balanced network, in that the true allocation consists of $M_{\mathrm{true}}$ communities with $50$ nodes. For the true probability matrix $Q$, we assume that each component is expressed by $q_{rs}=q+(p-q)I(r=s)$, where $q=0.1$, $p=0.8$ and $1\leq r\leq s\leq M_{\mathrm{true}}$. The assumption indicates that the edges are more easily drawn within the same community and less easily drawn between different communities. In the setup, we generate the $50$ data set.
We set $M-1|\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1)$, and employ (3.5) as a kernel and a prior, where $a_{Q}=1$ and $b_{Q}=1$. The values of the shape parameters, the evaluations and the MCMC setting are the same as those of Section 4.1.1.
Figure 4: Box plots of the posterior mean of the number of communities $M$. From left to right, shape parameters are set as $\gamma,\alpha=1,0.2,10^{-1},10^{-2},10^{-3}$.
We report the posterior mean of the number of communities in Figure 4. It is observed that the results are almost the same as those of Figure 2.
Table 7: Results of the rand index $\widehat{\mathrm{RI}}_{i}$ and posterior probability that the number of empty components is equal to zero ^$\widehat{\mathbb{P}}(M_{na}=0)$ (their standard deviations are shown in parentheses) averaged over 50 Monte Carlo replications.
$\widehat{\mathrm{RI}}_{i}$
$\alpha=\gamma=1$ $\alpha=\gamma=0.2$ $\alpha=\gamma=10^{-1}$ $\alpha=\gamma=10^{-2}$ $\alpha=\gamma=10^{-3}$
MFM-Inv-Ga $0.997$ $0.997$ $0.995$ $0.991$ $0.991$
$(0.010)$ $(0.019)$ $(0.032)$ $(0.044)$ $(0.044)$
MFM-Ga $0.995$ $0.996$ $0.980$ $0.987$ $0.910$
$(0.032)$ $(0.026)$ $(0.059)$ $(0.054)$ $(0.044)$
^$\widehat{\mathbb{P}}(M_{na}=0)$
$\alpha=\gamma=1$ $\alpha=\gamma=0.2$ $\alpha=\gamma=10^{-1}$ $\alpha=\gamma=10^{-2}$ $\alpha=\gamma=10^{-3}$
MFM-Inv-Ga $1.000$ $0.994$ $0.988$ $0.975$ $0.962$
$(0.001)$ $(0.009)$ $(0.029)$ $(0.044)$ $(0.094)$
MFM-Ga $0.960$ $0.500$ $0.311$ $0.089$ $0.068$
$(0.011)$ $(0.071)$ $(0.064)$ $(0.024)$ $(0.018)$
Table 7 also shows that the results of the posterior probability of $M_{na}$ are the same as Table 1. However, in terms of clustering accuracy for network data, MFM-Inv-Ga is higher and more accurate than MFM-Ga.
Table 8: Posterior probabilities and their standard deviations (shown in parentheses) of the number of components $M$ averaged over $50$ Monte Carlo replications. The highest value is bolded.
MFM-Inv-Ga
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
$\alpha=1$ $0.000$ $0.000$ $\bf{0.927}$ $0.073$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.243)$ $(0.243)$ $(0.000)$ $(0.000)$
$\alpha=0.2$ $0.000$ $0.012$ $\bf{0.966}$ $0.022$ $0.000$ $0.000$
$(0.000)$ $(0.084)$ $(0.111)$ $(0.073)$ $(0.001)$ $(0.000)$
$\alpha=10^{-1}$ $0.000$ $0.020$ $\bf{0.950}$ $0.027$ $0.002$ $0.000$
$(0.000)$ $(0.141)$ $(0.178)$ $(0.106)$ $(0.009)$ $(0.002)$
$\alpha=10^{-2}$ $0.000$ $0.036$ $\bf{0.936}$ $0.023$ $0.004$ $0.001$
$(0.000)$ $(0.179)$ $(0.184)$ $(0.035)$ $(0.008)$ $(0.004)$
$\alpha=10^{-3}$ $0.000$ $0.028$ $\bf{0.938}$ $0.023$ $0.007$ $0.004$
$(0.000)$ $(0.143)$ $(0.180)$ $(0.044)$ $(0.021)$ $(0.017)$
MFM-Ga
$\gamma=1$ $0.000$ $0.020$ $\bf{0.930}$ $0.049$ $0.002$ $0.000$
$(0.000)$ $(0.139)$ $(0.135)$ $(0.028)$ $(0.002)$ $(0.000)$
$\gamma=0.2$ $0.000$ $0.010$ $\bf{0.480}$ $0.307$ $0.133$ $0.070$
$(0.000)$ $(0.069)$ $(0.085)$ $(0.036)$ $(0.035)$ $(0.038)$
$\gamma=10^{-1}$ $0.000$ $0.035$ $\bf{0.300}$ $0.288$ $0.186$ $0.191$
$(0.000)$ $(0.104)$ $(0.059)$ $(0.042)$ $(0.040)$ $(0.070)$
$\gamma=10^{-2}$ $0.000$ $0.010$ $0.091$ $0.156$ $\bf{0.175}$ $0.566$
$(0.000)$ $(0.041)$ $(0.036)$ $(0.014)$ $(0.015)$ $(0.092)$
$\gamma=10^{-3}$ $0.000$ $0.006$ $0.069$ $0.131$ $\bf{0.160}$ $0.634$
$(0.000)$ $(0.028)$ $(0.023)$ $(0.017)$ $(0.012)$ $(0.085)$
From table 8, MFM-Inv-Ga has a much higher posterior probability at $M_{\text{true}}$ for all shape parameters than MFM-Ga. As a result, in the context of community detection, MFM-Inv-Ga also achieves better clustering and community estimation than MFM-Ga.
We also compared the computation time with Geng’s method (Geng et al., 2019), denoted MFM-Geng, when the shape parameters are $1.0$ and $10^{-3}$. The result in Table 9 indicates that MFM-Inv-Ga is on average more than twice as efficient as MFM-Geng in terms of computation time. The reason is why our algorithm does not update the label variable based on the restaurant process and MFM-Inv-Ga has a structure that is unlikely to produce empty components.
The MFM-Geng can estimate the number of clusters $k$, but cannot directly estimate the number of communities $M$. Furthermore, it is difficult in the MFM-Geng to estimate the hyperparameter of $q_{M}$, because it is necessary to perform complex series calculations, including its parameter. In summary, MFM-Inv-Ga (and MFM-Ga) are superior to MFM-Geng in terms of computation time, direct estimation of $M$, and estimation of an essential parameter.
Table 9: Comparison of CPU times in seconds (standard deviations are shown in parentheses) averaged over $50$ replications.
MFM-Inv-Ga MFM-Geng
$\alpha=\gamma=1$ $38.793$ $87.519$
$(2.894)$ $(3.719)$
$\alpha=\gamma=10^{-3}$ $38.361$ $83.423$
$(4.431)$ $(4.673)$
4.3.2 Dolphins social network data
The dolphins social network data is often used as a benchmark. Data can be obtained in http://www-personal.umich.edu/mejn/netdata/. The data is constructed as an undirected graph and expresses a small-scale animal social network with $64$ bottlenose dolphins off Doubtful Sound, New Zealand. Each node represents a dolphin, and an edge is drawn if two dolphins appear to be closely related to each other. In previous studies, it is well-known that the network has two communities. The details of the data are found in Lusseau et al. (2003).
As before, we compare the posterior distribution of the number of communities between MFM-Inv-Ga and MFM-Ga, and create a co-clustering matrix of MFM-Inv-Ga as quantifying uncertainty of clustering. In this analysis, we set $M-1\sim\text{Poisson}(1)$ and $a_{Q}=b_{Q}=3.0$.
Table 10: Posterior probabilities of the number of communities $M$ for dolphins social network data
MFM-Inv-Ga
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
$\alpha=1$ $0.000$ $0.997$ $0.003$ $0.000$ $0.000$ $0.000$
$\alpha=0.2$ $0.000$ $0.972$ $0.025$ $0.002$ $0.001$ $0.000$
$\alpha=10^{-1}$ $0.000$ $0.961$ $0.034$ $0.004$ $0.000$ $0.000$
$\alpha=10^{-2}$ $0.000$ $0.941$ $0.048$ $0.010$ $0.001$ $0.000$
$\alpha=10^{-3}$ $0.000$ $0.854$ $0.136$ $0.009$ $0.002$ $0.000$
MFM-Ga
$\gamma=1$ $0.000$ $0.949$ $0.049$ $0.002$ $0.000$ $0.000$
$\gamma=0.2$ $0.000$ $0.605$ $0.307$ $0.074$ $0.012$ $0.001$
$\gamma=10^{-1}$ $0.000$ $0.514$ $0.340$ $0.117$ $0.025$ $0.004$
$\gamma=10^{-2}$ $0.000$ $0.619$ $0.303$ $0.069$ $0.000$ $0.000$
$\gamma=10^{-3}$ $0.370$ $0.400$ $0.172$ $0.048$ $0.007$ $0.001$
From Table 10, the MFM-Inv-Ga is able to successfully estimate the number of communities $M=2$ regardless of the values of the shape parameter, while the MFM-Ga is not. We report a co-cluster matrix of MFM-Inv-Ga for $\alpha=1.0$ and MFM-Geng for $\gamma=1.0$, and the results are shown in Figure 5. The result of MFM-Inv-Ga is almost identical to the results reported in Geng et al. (2019). Furthermore, we confirmed that changing the value of $\alpha$ does not change the co-cluster matrices and the clustering solution based on MAP estimation. This shows the efficiency of the proposed method.
Figure 5: Co-clustering matrix with the MFM-Inv-Ga with $\alpha=1.0$ (left) and the MFM-Geng with $\gamma=1.0$ (right) for dolphins social network data.
5 Concluding remarks
We proposed a mixture of finite mixtures (MFM) model based on the normalized inverse Gaussian distribution and constructed an efficient posterior sampling algorithm based on Argiento and De Iorio (2022). The proposed method is a finite analog of the inverse Gaussian processes proposed by Lijoi et al. (2005). We illustrate the performance of the proposed method for clustering, density estimation, and community detection, compared to existing MFM models based on Dirichlet distribution (e.g., Miller and Harrison, 2018; Geng et al., 2019; Argiento and De Iorio, 2022). The proposed method is robust against the choice of hyper-parameter $\alpha$, and provided reasonable estimates of the number of components and communities compared to the MFM based on the Dirichlet prior distribution. Moreover, the proposed method also has a reasonable predictive performance in the sense of density estimation by suppressing the appearance of empty components.
The drawbacks of the proposed method are as follows. Some parameters involved in the model do not have closed-form marginal distributions because it is not easy to marginalize out $U_{n}$. For example, when we focus on clustering, obtaining the interpretable prior distribution for the number of clusters is very important (see e.g., Zito et al., 2023) to incorporate subjective prior information. However, the proposed model cannot lead to a tractable marginal prior distribution for the number of clusters. In the mixture of finite mixtures, the model consists of a distribution over a simplex based on the normalization of independent random variables. Therefore, the correlation between categories cannot be properly modeled. The normalized inverse Gaussian distribution has a negative covariance as well as the Dirichlet distribution. It may be inappropriate for data with positive correlations between categories, such as the proportion of symbiotic organisms present, disease complication data, or gene expression data. To address such a problem, it may be necessary to construct the MFM in a more general framework that removes the assumption of independence in the Norm-IFPP by Argiento and De Iorio (2022). The construction of MFM models using a more flexible distribution over a simplex that can also express positive correlations such as the generalized Dirichlet distribution (Wong, 1998) is an interesting future topic. Furthermore, the proposed model can be applied to spatial data (e.g., Geng et al., 2021) and functional data (e.g., Hu et al., 2023). For network data, it is expected to extend the MFM to network with weighted edges, degree-corrected stochastic block models, and mixed membership stochastic block models.
Acknowledgement
This work was supported by Japan Society for the Promotion of Science, the establishment of university fellowships towards the creation of science technology innovation Grant Number JPMJFS2129. This work is partially supported by the Japan Society for the Promotion of Science (grant number: 21K13835).

### v2

@@
 We consider a finite mixture model with an unknown number of components:
-$f(y|\theta_{1},\dots,\theta_{M},\pi)=\sum_{j=1}^{M}\pi_{j}f(y|\theta_{j}),$ (1.1)
+$f(y\mid\theta_{1},\dots,\theta_{M},\pi)=\sum_{j=1}^{M}\pi_{j}f(y\mid\theta_{j}),$ (1.1)
 where $M\in\{1,2,\dots\}$ is the number of components, $y\in\mathbb{R}^{d}$ is a $d$-dimensional observation vector, and $\pi=(\pi_{1},\dots,\pi_{M})^{\top}$ is the mixing weight vector, satisfying $\sum_{j=1}^{M}\pi_{j}=1$ and $\pi_{j}>0$.
-For $j=1,\dots,M$, $f(\cdot|\theta_{j})$ is a probability density or probability mass function parameterized by the component-specific parameter $\theta_{j}$.
-The function $f(\cdot|\theta_{j})$ is also called the kernel of the mixture model.
+For $j=1,\dots,M$, $f(\cdot\mid\theta_{j})$ is a probability density or probability mass function parameterized by the component-specific parameter $\theta_{j}$.
+The function $f(\cdot\mid\theta_{j})$ is also called the kernel of the mixture model.
 Each observation is assumed to arise from one of the $M$ components, and each component is weighted based on its frequency of occurrence.
 Mixture models are important statistical models used in model-based clustering and density estimation because they can model complex data-generating distributions in random phenomena by using multiple components.
-Finite mixture models have been applied in many fields, for instance, sociology (Handcock et al., 2007), economics (Frühwirth-Schnatter and Kaufmann, 2008) and genetics (McLachlan et al., 2002) and so on.
-In application, the determination of $M$ is a very important issue.
+Finite mixture models have been applied in many fields, for instance, sociology (handcock2007model), economics (frohwirth2008model) and genetics (mclachlan2002mixture) and so on.
+In applications, the determination of $M$ is a very important issue.
 Correctly estimating $M$ improves model interpretability, the accuracy of kernel parameter estimates and model predictions, and reduces computation time.
-Although various methods have been proposed for the selection of $M$, such as model selection criteria and hypothesis testing, using these criterion makes it difficult to quantify the uncertainty of $M$, and also introduces the bias that model selection is carried out.
-The comprehensive review of (finite) mixture models is provided by McLachlan (2000); Frühwirth-Schnatter (2006); Frühwirth-Schnatter et al.
-(2019); McLachlan et al.
-(2019).
+Although various methods have been proposed for the selection of $M$, such as model selection criteria and hypothesis testing, using these criteria make it difficult to quantify the uncertainty of $M$, and also introduces the bias that model selection is carried out.
+A comprehensive review of (finite) mixture models is provided by mclachlan2000finite; fruhwirth2006finite; fruhwirth2019handbook; mclachlan2019finite.
 In this paper, we focus on the case where $M$ is finite and make a clear distinction between the number of components $M$ and the number of clusters $k$.
-In general, the number of components is written as $M=k+M_{na}$, where $k$ is the number of components for which data are actually assigned, and $M_{na}$ is the number of empty components (see also Argiento and De Iorio, 2022).
+In general, the number of components is written as $M=k+M_{na}$, where $k$ is the number of components for which data are actually assigned, and $M_{na}$ is the number of empty components (see also argiento2022infinity).
+Although we also estimate $M$ in this paper, there are two important issues associated with its inference.
+First, for any finite data set, $M$ is not a well-defined quantity and cannot be directly inferred from the data.
+In certain application domains of mixture models, such as species sampling in biology (see e.g., argiento2014density), it is even expected that data from some components may not be observed at all.
+Thus, given the data, we can only make predictions about $M$, and the prior on $M$ and the prior on the weight distribution play a crucial role.
+Second, the posterior distribution of $M$ is highly sensitive to model misspecification.
+Most consistency results for $M$ are derived under the assumption of model correctness (nobile1994bayesian; miller2023consistency).
+Therefore, when attempting to infer the number of mixture components, it is critically important to ensure that the data-generating process is appropriately captured.
 In Bayesian analysis, a finite mixture model that assumes prior distributions for the mixing weights and the number of components is usually employed (see, e.g.
-Nobile, 1994; Miller and Harrison, 2018).
+nobile1994bayesian; miller2018mixture).
 Such a model is also called a mixture of finite mixtures (MFM), and the model is often used, as well as infinite mixture models represented by Dirichlet process mixture models.
-Although the reversible jump algorithm (Green, 1995; Richardson and Green, 1997) has been used to obtain samples from the posterior distribution based on MFM, it faces significant computational and implementation challenges.
+Although the reversible jump algorithm (green1995reversible; richardson1997bayesian) has been used to obtain samples from the posterior distribution based on MFM, it faces significant computational and implementation challenges.
 Methods based on marginal likelihoods have also been proposed (see, e.g.
-Nobile and Fearnside, 2007), while, as with reversible jump, the computational aspect is an issue.
-A method called sparse MFM (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016) has also been proposed, which focuses on estimating the number of clusters $k$ by using the overfitting model with a large fixed $M$, and choosing the prior distribution of the component rate well.
-Under same conditions, they showed that $k$ has consistency for true $M$.
-Since it allows for many empty components, we cannot estimate the number of components $M$ and it is not easy to quantify the uncertainty.
+nobile2007bayesian), while, as with reversible jump, the computational aspect is an issue.
 Recently, the use of nonparametric Bayesian methods for MFM to estimate $M$ has attracted much attention.
-In Miller and Harrison (2018), the exchangeable partition distribution is derived by marginalizing out $M$, and the restaurant process is constructed to overcome computational difficulties.
-In Frühwirth-Schnatter et al.
-(2021), a generalized MFM in which the parameters of the Dirichlet distribution depend on $M$ is proposed.
-Miller and Harrison (2018) derived the theoretical result that the marginal posterior probabilities of $k$ and $M$ are asymptotically equivalent when the sample size is large, while only a sampling of $k$ is obtained as with general nonparametric Bayesian methods.
-Thus, the number of components $M$ cannot be sampled directly from the posterior distribution.
-Although Frühwirth-Schnatter et al.
-(2021)’s method can sample $M$ directly, as a sparse MFM, it has the disadvantage of producing a large number of empty components and overestimating $M$.
-In addition, most MFM studies use a symmetric Dirichlet distribution as the mixing weight vector, and the choice of a hyperparameter is very important.
-Although it is possible to set a prior distribution and learn from the data, the Metropolis-Hastings algorithm is required.
-The main purposes of this paper are 1) sampling $M$ directly and efficiently compared to MFM based on the Dirichlet distribution; 2) estimating $M$ reasonably by suppressing $M_{na}$; 3) proposing a MFM that is robust to the choice of hyperparameters of the distribution on a simplex.
-To this end, we employ a normalized inverse Gaussian distribution as the mixing weight vector of the finite mixture model.
+In miller2018mixture, the exchangeable partition distribution is derived by marginalizing out $M$, and the restaurant process is constructed to overcome computational difficulties.
+Note that the result of Miller and Harrison (2018) has some connection with Pitman-Yor process prior (see e.g.
+Section 4.1 in fruhwirth2021generalized.
+In fruhwirth2021generalized, a generalized MFM in which the parameters of the Dirichlet distribution depend on $M$ is proposed.
+They showed that this model class may be regarded as a Bayesian nonparametric mixture outside the class of Gibbs-type priors, and proposed a novel sampling algorithm so-called telescoping sampling.
+In argiento2022infinity, they developed an MFM framework based on a nonparametric Bayesian prior, the normalized independent finite point process.
+By introducing a gamma latent variable, their approach encompasses a broad class of distributions on the simplex and yields an efficient blocked Gibbs sampling derived from the posterior random measure.
+Most studies on MFMs employ a (symmetric) Dirichlet distribution for the mixing weights because of its analytical convenience.
+argiento2022infinity primarily consider not only the Dirichlet distribution but also a family of simplex distributions constructed from $\sigma$-stable laws.
+However, the $\sigma$-stable example is treated as a central case, since their paper emphasizes that a restaurant process based computational algorithm can be constructed whenever the Laplace transform of the weight distribution is available.
+In this paper, we study a new MFM based on the normalized inverse Gaussian distribution, building on the results of argiento2022infinity.
+We compare the proposed MFM with the MFM on the Dirichlet prior, focusing on the distributional differences from the Dirichlet case.
 To the best of our knowledge, there are few studies using the normalized inverse Gaussian distributions as the mixing weights in the field of finite mixture models.
-In Bayesian nonparametric inference with infinite mixtures, Lijoi et al.
-(2005) proposed a normalized inverse Gaussian process and showed some analytical results of the posterior distribution under the process.
-Similarly, since Miller and Harrison (2018) is a finite version of the Dirichlet process, the model presented in this paper corresponds to a kind of finite alternative of Lijoi et al.
-(2005).
-Furthermore, leveraging data augmentation with a latent gamma random variable and the result of Argiento and De Iorio (2022), we construct an efficient posterior sampling algorithm.
+In Bayesian nonparametric inference with infinite mixtures, lijoi2005hierarchical proposed a normalized inverse Gaussian process and showed some analytical results of the posterior distribution under the process.
+Furthermore, we clarify the relationship between the blocked Gibbs sampling and the telescoping sampling, and extend the telescoping sampling so that it can be applied to general distributions on the simplex.
 This paper is organized as follows.
-In Section 2, we introduce the MFM and its equivalent representation using discrete probability measures.
-We also discuss the data augmentation and the conditional posterior distribution of $M_{na}$, which are crucial for constructing the computational algorithm.In Section 3, we present the proposal method as well as the posterior computation algorithm.
+In Section 2, we provide a brief review of MFMs and its equivalent representation using discrete probability measures.
+We also discuss the data augmentation and the conditional posterior distribution of $M_{na}$, which are crucial for constructing the computational algorithm.
+In Section 3, we present the proposal method as well as the posterior computation algorithm.
 In Section 4, we illustrate some numerical studies to compare the proposed method with existing methods.
@@
 2 Mixture of finite mixtures
-We introduce a mixture of finite mixtures and its equivalent representation of the discrete probability measure.
+We present a brief overview of a MFM and its equivalent representation of the discrete probability measure.
 The proposed model and the posterior computation algorithm presented in the next section are largely based on the basic model presented in this section.
 2.1 Formulation
-Let each observations $Y_{1},\dots,Y_{n}$ be univariate or multivariate in an Euclidean space. A MFM model is a statistical model defined by the following hierarchical representation.
-$\begin{split}&M-1\sim q_{M},\quad q_{M}\ \text{is a probability mass function (p.m.f.) on}\ \{0,1,2,\dots\},\\ &\pi|M\sim P_{\pi}(\pi|M),\\ &c_{i}|M,\pi\sim\mathrm{Categorical}_{M}(\pi),\\ &\xi_{m}|M\overset{\mathrm{i.i.d}}{\sim}p_{0}(\xi),\quad m=1,\dots,M,\\ &\theta_{i}|c_{i},\xi\overset{\mathrm{ind}}{\sim}\delta_{\xi_{c_{i}}}(\theta_{i}),\quad i=1,\dots n,\\ &Y_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}|\theta_{i}),\quad i=1,\dots,n,\end{split}$ (2.1)
-where $f(y|\theta_{i})$ is a parametric model with parameter $\theta_{i}$, $\delta_{x}(\cdot)$ is a point mass at $x$ and $\xi$ is a element of a vector $\xi=(\xi_{1},\dots,\xi_{M})^{\top}$.
-$p_{0}(\xi)$ is a prior density function of the parameter in mixture components.
+Let each observation $Y_{1},\dots,Y_{n}$ be univariate or multivariate in a Euclidean space. An MFM is a statistical model defined by the following hierarchical representation.
+$\begin{split}&M-1\sim q_{M},\quad q_{M}\ \text{is a probability mass function (p.m.f.) on}\ \{0,1,2,\dots\},\\ &\pi\mid M\sim P_{\pi}(\pi\mid M),\\ &c_{i}\mid M,\pi\sim\mathrm{Categorical}_{M}(\pi),\quad i=1,\dots,n,\\ &\tau_{m}\mid M\overset{\mathrm{i.i.d}}{\sim}p_{0}(\tau),\quad m=1,\dots,M,\\ &\theta_{i}\mid c_{i},\tau_{1:M}\overset{\mathrm{ind}}{\sim}\delta_{\tau_{c_{i}}}(\theta_{i}),\quad i=1,\dots n,\\ &Y_{i}\mid\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}\mid\theta_{i}),\quad i=1,\dots,n,\end{split}$ (2.1)
+where $f(y\mid\theta_{i})$ is a parametric model with parameter $\theta_{i}$, $\delta_{x}(\cdot)$ is a point mass at $x$ and $\tau_{m}$ is an element of a vector $\tau_{1:M}=(\tau_{1},\dots,\tau_{M})^{\top}$.
+$p_{0}(\tau_{1:M})$ is a prior density function of the parameter in mixture components.
 The parameter space is denoted by $\Theta\subset\mathbb{R}^{d}$.
 Each $c_{i}$ is a latent allocation indicates to which component each observation is allocated.
-Given the number of components $M$, $P_{\pi}(\pi|M)$ is a probability distribution on the $M$-dimensional unit simplex, in that, this is the prior distribution for the mixing weights.
-We focus the case where the mixing weights can be further hierarchically expressed as follows:
-$\begin{split}&S_{m}|M\overset{\mathrm{ind}}{\sim}h,\quad h\ \text{is a probability density function (p.d.f.) on}\ (0,\infty),\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi|M),\quad T=\sum_{m=1}^{M}S_{m},\end{split}$ (2.2)
-where $S_{1},\dots,S_{M}$ is unnormalized weights and $\pi$ is (normalized) mixing weight vector.
+Given the number of components $M$, $P_{\pi}(\pi\mid M)$ is a probability distribution on the $M$-dimensional unit simplex, in that, this is the prior distribution for the mixing weights.
+Let $h$ be a probability density function (p.d.f) on $\mathbb{R}_{+}$. We focus on the case where the mixing weights can be further hierarchically expressed as follows:
+$\begin{split}&S_{m}\mid M\overset{\mathrm{ind}}{\sim}h,\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi\mid M),\quad T=\sum_{m=1}^{M}S_{m},\end{split}$ (2.2)
+where $S_{1},\dots,S_{M}$ are unnormalized weights and $\pi$ is (normalized) mixing weight vector.
 This representation is the most basic way to generate a probability distribution on the $d$-dimensional unit simplex $\mathbb{S}_{d}:=\{w=(w_{1},\dots,w_{d});w_{i}\geq 0,\sum_{i=1}^{d}w_{i}=1\}$.
 One of the most famous examples of a distribution on $\mathbb{S}_{d}$ is the Dirichlet distribution.
-If $S_{m}|M\sim\text{Gamma}(\gamma_{i},1)$ in (2.2), $\pi\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$, where $\text{Gamma}(a,b)$ is the gamma distribution with shape parameter $a>0$ and scale parameter $\beta>0$, and $\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$ is the Dirichlet distribution with parameter $\gamma=(\gamma_{1},\dots,\gamma_{d})>0$.
-If $\gamma_{1}=\cdots=\gamma_{d}=:\gamma$, the distribution obtained by normalization is $\text{Dirichlet}(\gamma,\dots,\gamma)$.
-The symmetric Dirichlet distribution is often used in the framework of MFM, because of conjugacy with categorical distributions and simplicity of computation.
-The symmetric structure of $\text{Dirichlet}(\gamma,\dots,\gamma)$ is also essential for marginalizing out $M$ and deriving the exchangeable partition distribution.
+If $S_{i}\sim\text{Gamma}(\gamma_{i},1)$ in (2.2), $\pi\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$, where $\text{Gamma}(a,b)$ is the gamma distribution with shape parameter $a>0$ and rate parameter $b>0$, and $\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$ is the Dirichlet distribution with parameter $(\gamma_{1},\dots,\gamma_{d})$.
+When we set $\gamma_{1}=\cdots=\gamma_{d}\ (=\gamma>0)$, we have the symmetric Dirichlet distribution $\text{Dirichlet}(\gamma,\dots,\gamma)$, and this distribution is often used as a prior distribution for mixing weight vector in the framework of MFM, because of conjugacy with ctegorical distributions and simplicity of computation.
+The symmetric structure of $\text{Dirichlet}(\gamma,\dots,\gamma)$ is also essential for marginalizing out $M$ and deriving the exchangeable partition probability function (EPPF).
+Also, if the hyperparameters of $h$ depend on $M$ in (2.2), the model is referred to as a generalized MFM:
+$\begin{split}&S_{m}\mid M\overset{\mathrm{ind}}{\sim}h_{M},\quad m=1,\dots,M,\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi\mid M),\quad T=\sum_{m=1}^{M}S_{m}.\end{split}$ (2.3)
+In particular, if the shape parameters are divided by $M$, the model is referred to as a dynamic MFM (DMFM).
+The Dirichlet case, where $h_{M}$ is $\text{Gamma}(\gamma/M,1)$, is discussed in fruhwirth2021generalized.
+On the other hand, fruhwirth2021generalized referred to MFM defined by (2.1), in which the distributional parameters of $S_{m}$ do not depend on $M$, as static MFM.
+In this paper, we simply refer to such static MFM as MFM.
 However, it is known that the estimation result is sensitive to the choice of the shape parameter $\gamma$.
-For example, Miller and Harrison (2018) recommended to use $\gamma=1$ as a default choice and the value works well in many cases.
-The use of a small $\gamma$ was recommended in terms of sparse or generalized MFMs (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016; Frühwirth-Schnatter et al., 2021).
-A small value of $\gamma$ indicates that many empty components can be created.
-Although most previous studies allow for the appearance of the empty components, there are several problems.
-First, $M$ tends to be overestimated, making the interpretation of the model difficult.
-Second, the existing of the empty components decreases the predictive performance of the model, as we will see in a later section through density estimations.
-Finally, they lead to an increase in computation time.
-When we focus on the estimation of the number of components $M$, it is desirable to have few empty components.
-In other words, the discrepancy between $M$ and $k$ should be small.
-To achieve this, we use the normalized inverse Gaussian distribution as mixing weights, instead of the Dirichlet distribution.
+For example, miller2018mixture recommended to use $\gamma=1$ as a default choice and the value works well in many cases.
+On the other hand, in the sparse MFM, dividing by a large $M$ effectively corresponds to using a small value of $\gamma$.
+Moreover, because $\gamma$ strongly influences both the prior distribution of the number of clusters and the induced partition distribution, it is necessary to assign an appropriate prior to $\gamma$.
+For further discussion of broader prior specifications, see greve2022spying.
 2.2 Equivalent representations using discrete probability measures
-We give an equivalent representation of the MFM.
-We can construct a discrete measure $P(\cdot)=\sum_{m=1}^{M}\pi_{m}\delta_{\xi_{m}}(\cdot)$ in the parameter space $\Theta$, almost surely, where $M$, $\xi$ and $\pi$ are realizations from the distributions $q_{M}$, $p_{0}$ and $P_{\pi}$, respectively.
+Following argiento2022infinity, we give an equivalent representation of the MFM.
+We can construct a discrete measure $P(\cdot)=\sum_{m=1}^{M}\pi_{m}\delta_{\tau_{m}}(\cdot)$ in the parameter space $\Theta$, almost surely, where $M$, $\tau$ and $\pi$ are realizations from the distributions $q_{M}$, $p_{0}$ and $P_{\pi}$, respectively.
 Let $\theta_{1},\dots,\theta_{n}$ be random variables according to $P$.
 Then, the model (2.1) is equivalent to the following hierarchical representation using a random measure $P$.
-$\begin{split}&Y_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}|\theta_{i}),\quad i=1,\dots,n,\\ &\theta_{1},\dots,\theta_{n}|P\overset{\mathrm{i.i.d}}{\sim}P,\\ &P\sim\mathcal{P}(q_{M},h,p_{0}),\end{split}$ (2.3)
+$\begin{split}&Y_{i}\mid\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}\mid\theta_{i}),\quad i=1,\dots,n,\\ &\theta_{1},\dots,\theta_{n}\mid P\overset{\mathrm{i.i.d}}{\sim}P,\\ &P\sim\mathcal{P}(q_{M},h,p_{0}),\end{split}$ (2.4)
 where $\mathcal{P}$ is a probability distribution of $P$ with parameters $q_{M}$, $h$ and $p_{0}$.
-The representation (2.3) is the MFM described as a nonparametric Bayesian framework with infinite mixtures.
-If we replace $\mathcal{P}$ with the Dirichlet process, (2.3) represents the well-known Dirichlet process mixture models (e.g., Escobar and West, 1995).
-If we replace $\mathcal{P}$ with the normalized inverse Gaussian process, (2.3) represents the normalized inverse Gaussian process mixture models (Lijoi et al., 2005).
-For MFM, Argiento and De Iorio (2022) proposed a normalized independent finite point process (Norm-IFPP), which is a class of flexible prior distributions for $P$.
+The representation (2.4) is the MFM described in the same framework as infinite mixtures.
+If we replace $\mathcal{P}$ with the Dirichlet process, (2.4) represents the well-known Dirichlet process mixture models (e.g., escobar1995bayesian).
+If we replace $\mathcal{P}$ with the normalized inverse Gaussian process, (2.4) represents the normalized inverse Gaussian process mixture models (lijoi2005hierarchical).
+For MFMs, argiento2022infinity proposed a normalized independent finite point process (Norm-IFPP), which is a class of flexible prior distributions for $P$.
 We employ the representation (2.4) with the Norm-IFPP.
@@
 We can directly estimate $M$ and $k$.
-This is a major difference from Miller and Harrison (2018), which estimates $M$ indirectly through the number of clusters $k$.
 Furthermore, an efficient Gibbs sampler can be constructed by incorporating the data augmentation with a latent Gamma random variable.
@@
 This is the key to building an efficient MCMC algorithm for the proposed model.
-Our study is in the spirit of nonparametric Bayes in that it accounts for uncertainty in the prior distribution of the parameters by means of a random measure based on the Norm-IFPP.
-The details of Norm-IFPP and and the independent finite point process (IFPP) are given in Argiento and De Iorio (2022).
+The details of Norm-IFPP and the independent finite point process (IFPP) are given in argiento2022infinity.
 2.3 Data augmentation and conditional distribution of $M_{na}$
@@
 We introduce the data augmentation in models (2.1) and (2.2).
-This technique is employed in James et al.
-(2009) and Argiento and De Iorio (2022).
+This technique is employed in james2009posterior and argiento2022infinity.
 Let $\mathcal{M}_{a}$ and $\mathcal{M}_{na}$ be the allocated and unallocated index sets, respectively.
-The conditional joint distribution of $\xi=(\xi_{1},\dots,\xi_{M})^{\top}$ and $S=(S_{1},\dots,S_{M})^{\top}$ given $M$ and a label vector $c$ is
-$p(S,\tau|M,c)$ $\propto\left(\prod_{i=1}^{n}\frac{S_{c_{i}}}{T}\right)\left(\prod_{m=1}^{M}h(s_{m})p(\tau_{m})\right)$
-$=\left(\prod_{m\in\mathcal{M}^{a}}\left(\frac{S_{m}}{T}\right)^{n_{m}}h(s_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}^{na}}h(s_{m})p(\tau_{m})\right),$
-where $n_{m}=\#\{i:c_{i}=m\}$.
-Since this equation consists of categorical distributions, conjugacy with such a distribution is required as $P_{\pi}$ for Gibbs sampling.
-This restriction is relaxed by using latent a random variable $U_{n}|T\sim\mathrm{Gamma}(n,T)$, where $T=\sum_{m=1}^{M}S_{m}$.
-In fact,
-$p(S,\tau,U_{n}|M,c)\propto u^{n-1}\left(\prod_{m\in\mathcal{M}}e^{-S_{m}u}S_{m}^{n_{m}}h(s_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}^{na}}e^{-S_{m}u}h(s_{m})p(\tau_{m})\right).$
+The conditional joint distribution of $\tau=(\tau_{1},\dots,\tau_{M})^{\top}$ and $S=(S_{1},\dots,S_{M})^{\top}$ given $M$ and a label vector $c$ is
+$p(S,\tau\mid M,c)$ $\propto\left(\prod_{i=1}^{n}\frac{S_{c_{i}}}{T}\right)\left(\prod_{m=1}^{M}h(S_{m})p(\tau_{m})\right)$
+$=\left(\prod_{m\in\mathcal{M}_{a}}\left(\frac{S_{m}}{T}\right)^{n_{m}}h(S_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}_{na}}h(S_{m})p(\tau_{m})\right),$
+where $n_{m}=\#\{i:c_{i}=m\}$ and $T=\sum_{m=1}^{M}S_{m}$, which depends on $S$. Furthermore, using $U_{n}\mid T\sim\mathrm{Gamma}(n,T)$, we obtain
+$p(S,\tau,U_{n}\mid M,c)\propto u^{n-1}\left(\prod_{m\in\mathcal{M}_{a}}e^{-S_{m}U_{n}}S_{m}^{n_{m}}h(S_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}_{na}}e^{-S_{m}U_{n}}h(S_{m})p(\tau_{m})\right).$
 Thus, the conditional distribution of each $S_{m}$ is
-$S_{m}|U_{n},M,c\overset{\mathrm{ind}}{\sim}e^{-S_{m}u}S_{m}^{n_{m}}h(s_{m}),\quad m\in\mathcal{M}_{a},$ (2.4)
-$S_{m}|U_{n},M,c\overset{\mathrm{i.i.d}}{\sim}e^{-S_{m}u}h(s_{m}),\quad m\in\mathcal{M}_{na}.$ (2.5)
+$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{ind}}{\sim}e^{-S_{m}u}S_{m}^{n_{m}}h(S_{m}),\quad m\in\mathcal{M}_{a},$ (2.5)
+$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{i.i.d}}{\sim}e^{-S_{m}u}h(S_{m}),\quad m\in\mathcal{M}_{na}.$ (2.6)
 If it is easy to generate random variables from the distributions in (2.5) and (2.6), an efficient Gibbs sampling algorithm can be constructed.
@@
 Thus, the selection of $h$ is essential in MCMC updates.
-Under this data augmentation, the conditional distribution of $M_{na}$ is established in Theorem 5.1 in Argiento and De Iorio (2022).
+Under this data augmentation, the conditional distribution of $M_{na}$ is established in Theorem 5.1 in argiento2022infinity.
 This theorem states that if $\mathcal{P}(q_{M},h,p_{0})$ follows a Norm-IFPP, then the posterior distribution of the random measure $P$, given $U_{n}$, is a superposition of a finite point process with fixed points and an IFPP.
@@
 The conditional distribution of $M_{na}$ is given by
-$\mathbb{P}(M_{na}=m|\theta,U_{n}=u)\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=1,2,\dots,$ (2.6)
-where $\theta=(\theta_{1},\dots,\theta_{n})^{\top}$, $k$ is the number of cluster (unique values of $\theta$) and $\psi$ is a Laplace transform of $h$.
+$\mathbb{P}(M_{na}=m\mid\theta,U_{n}=u,k)\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=1,2,\dots,$ (2.7)
+where $\theta=(\theta_{1},\dots,\theta_{n})^{\top}$, $k$ is the number of clusters (unique values of $\theta$) and $\psi$ is the Laplace transform of $h$.
 We sample $M_{na}$ from (2.7) in the MCMC algorithm.
 Then, we straightforwardly obtain $M$ by adding $k$ to $M_{na}$.
-Remark 2.1.
-Although we use the data augmentation $U_{n}|T\sim\text{Gamma}(n,T)$, the distributions (2.4) and (2.5) are generally complex.
-In addition, the result of Argiento and De Iorio (2022) applies only to $h$ that have a Laplace transform.
-Thus, the inverse Gaussian distribution is one of the few examples that satisfy both computational and theoretical constraints.
 3 Methodology
-In this section, we propose a mixture of finite mixtures with normalized inverse Gaussian weights. Moreover, we develop an efficient posterior sampling algorithm based on Argiento and De Iorio (2022).
+In this section, we propose a mixture of finite mixtures with normalized inverse Gaussian weights.
+The posterior inference based on the proposed model can be efficiently implemented using the blocked Gibbs sampling proposed by argiento2022infinity.
 3.1 Mixture of finite mixtures with normalized inverse Gaussian weights
-We propose a mixture of finite mixtures with normalized inverse Gaussian weights (denoted by MFM-Inv-Ga), where the notation explicitly specifies $h$ because it is essential for computation.
+We propose a mixture of finite mixtures with normalized inverse Gaussian weights (denoted by MFM-IGau), where the notation explicitly specifies $h$ because it is essential for computation.
 Our proposal model only requires (2.2) to be
-$S_{m}|M\overset{\mathrm{ind}}{\sim}\text{Inv-Ga}(\alpha,1),\quad\alpha>0,$
-$\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim\text{Norm-Inv-Ga}(\alpha,\dots,\alpha),$
-[…]
-$h(s)=\frac{\alpha}{\sqrt{2\pi}}s^{-3/2}\exp\left(-\frac{1}{2}\left(\frac{\alpha^{2}}{s}+\beta^{2}s\right)+\beta\alpha\right),$ (3.1)
+$S_{m}\mid M\overset{\mathrm{ind}}{\sim}\text{IGau}(\alpha,1),\quad\alpha>0,$
+$\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim\text{NIGau}(\alpha,\dots,\alpha),$
+where $\text{IGau}(\alpha,\beta)$ is the inverse Gaussian distribution with shape parameter $\alpha>0$ and scale parameter $\beta>0$, and $\text{NIGau}(\alpha_{1},\dots,\alpha_{d})$ is the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0$ for $i=1,\dots,d$.
+The probability density function of a random variable $S$ according to $\text{IGau}(\alpha,\beta)$ is given by
+$h(s)=\frac{\alpha}{\sqrt{2\pi}}s^{-3/2}\exp\left(-\frac{1}{2}\left(\frac{\alpha^{2}}{s}+\beta^{2}s\right)+\beta\alpha\right),\quad s>0,$ (3.1)
 where $\alpha>0$ is the shape parameter and $\beta>0$ is the scale parameter. The mean and variance of $S$ are given by
 $\mathbb{E}[S]=\frac{\alpha}{\beta},\quad\mathrm{Var}\left[S\right]=\frac{\alpha}{\beta^{3}}.$ (3.2)
-From Lijoi et al. (2005), the probability density function of the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0,\ i=1,\dots,d$ is given by
+Let $\psi_{\text{IGau}}(u)$ be the Laplace transforms of $\text{IGau}(\alpha,1)$. $\psi_{\text{IGau}}(u)$ is the following:
+$\psi_{\text{IGau}}(u)=\exp\left(\alpha\left(1-\sqrt{1+2u}\right)\right),\quad u\geq 0.$ (3.3)
+Also, cumulant $\kappa_{\mathrm{IGau}}(n,u):=(-1)^{n}\frac{d^{n}}{du^{n}}\psi_{\mathrm{IGau}}(u)$ is
+$\kappa_{\mathrm{IGau}}(u;n)=\alpha^{n}\psi_{\text{IGau}}(u)(1+2u)^{-n/2}\frac{K_{n-1/2}\left(\alpha\sqrt{1+2u}\right)}{K_{1/2}\left(\alpha\sqrt{1+2u}\right)},$ (3.4)
+where $K_{m}$ is the modified Bessel functions of the second kind of order $m$.
+The derivation of (3.4) is provided in the Supplementary Material.
+From lijoi2005hierarchical, the probability density function of the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0,\ i=1,\dots,d$ is given by
 $\begin{split}f(\pi)=&\frac{e^{\sum_{i=1}^{d}\alpha_{i}}\prod_{i=1}^{d}\alpha_{i}}{2^{d/2-1}\pi^{d/2}}\times K_{-d/2}\left(\sqrt{\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})}\right)\\ &\times\left(\pi_{1}^{3/2}\cdots\pi_{d-1}^{3/2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{3/2}\times\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})^{d/4}\right)^{-1},\end{split}$ (3.5)
-where $\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})=\sum_{i=1}^{d-1}\alpha_{i}^{2}/\pi_{i}+\alpha_{d}^{2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{-1}$ and $K_{m}$ is the modified Bessel functions of the third kind of order $m$ (see, also Ghosal and van der Vaart, 2017).
+where $\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})=\sum_{i=1}^{d-1}\alpha_{i}^{2}/\pi_{i}+\alpha_{d}^{2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{-1}$.
+The normalized inverse Gaussian distribution is less informative and more robust to parameter specification than the Dirichlet distribution.
+In the context of DMFM, shape parameters such as $\gamma/M$ or $\alpha/M$ become small, and information measures such as entropy and the Gini index of $\mathrm{Dirichlet}(\gamma/M,\dots,\gamma/M)$ also tends to be very small.
+Therefore, it is desirable to incorporate learning for $\gamma$.
+In contrast, those of $\mathrm{NIGau}(\alpha/M,\dots,\alpha/M)$ does not decrease as easily, so stable posterior inference can be achieved regardless of the choice of $\alpha$.
+This fact is confirmed through the numerical experiment in Section 4.1.1.
+Further analysis of the Gini index is provided in the Supplemental Material.
+Also, in the context of infinite mixtures, it is known that the prior on the number of clusters induced by the normalized inverse Gaussian process has heavier tails than that induced by the Dirichlet process (lijoi2005hierarchical).
+It is expected that this robustness carries over to the finite mixture setting as well.
+These provide the main motivations for using the normalized inverse Gaussian distribution as mixing weights.
 Due to the breakdown of conjugacy with the categorical distribution, constructing an efficient Gibbs sampler is challenging.
-Because of the complexity of (3.3), the calculation in Miller and Harrison (2018) is intractable: constructing an MCMC algorithm based on the restaurant process by marginalizing out $M$ is extremely challenging.
-To overcome this difficulty, we use the data augmentation in Subsection 2.3 and the method proposed in Argiento and De Iorio (2022).
-As a result, it is sufficient to use not (3.3) but (3.1) and (3.8) to construct our MCMC algorithm.
+Because of the complexity of (3.5), the calculation in miller2018mixture is intractable: constructing MCMC algorithm based on the restaurant process by marginalizing out $M$ is extremely challenging.
+To overcome this difficulty, we use the data augmentation technique and the sampling method proposed by argiento2022infinity.
+As a result, it is no longer necessary to deal directly with the normalized inverse Gaussian distribution (3.5), and the simple MCMC algorithm can be constructed using only the inverse Gaussian distribution (3.1) and its Laplace transform.
 3.2 Posterior computation
-We present a fast and efficient posterior computation algorithm for the proposed method.
-To this end, we adopt the blocked Gibbs sampling scheme in Argiento and De Iorio (2022).
-Let $\eta$ be a hyper-parameter of the prior distribution $q_{M}$, and $p(\cdot)$ be a joint prior density function except for $h$ and $q_{M}$.
+In this section, we present a fast and efficient posterior computation algorithm for the proposed method.
+To this end, we adopt the blocked Gibbs sampling scheme in argiento2022infinity.
+Moreover, we discuss the blocked Gibbs sampling in detail and explain the computational advantages of using inverse Gaussian weights.
+Let $\eta$ and $\rho$ be hyperparameters of $q_{M}$ and $h$, respectively, and $p(\cdot)$ be a joint prior density function except for $h$ and $q_{M}$.
 $n_{m}=\#\{i;c_{i}=m\}$ is the size of the $m$th cluster.
 $\psi(u)$ is the Laplace transform of $h$, which is defined by $\psi(u):=\mathbb{E}[e^{-uS_{m}}]$ for $u\geq 0$.
-Algorithm 1 Block Gibbs sampler
+$\kappa(u;n)$ is the cumulant of $h$, which is defined by $\kappa(u;n):=(-1)^{n}\frac{d^{n}}{du^{n}}\psi(u)$.
+Note that $\psi(u)$ and $\kappa(u;n)$ depend on $\rho$.
+Algorithm 1 blocked Gibbs sampling (argiento2022infinity)
 Step 0. Set initial values.
 Step 1. Sample $U_{n}$ from $\mathrm{Gamma}(n,T)$, where $T=\sum_{m=1}^{M}S_{m}$.
-  Step 2. Sample $c_{i},for\ i=1,\dots,n,$ the following discrete probability distribution defined by each probabilities
-$\mathbb{P}(c_{i}=j|\mathrm{rest})\propto S_{j}f(y_{i}|\tau_{j}),\quad j=1,\dots,M.$
+Step 2. Sample $c_{i}$, for $i=1,\dots,n$ from the following discrete probability distribution defined by each probability
+$\mathbb{P}(c_{i}=j\mid\mathrm{rest})\propto S_{j}f(y_{i}\mid\tau_{j}),\quad j=1,\dots,M.$
 Step 3. Sample the hyperparameter $\eta$ of the $q_{M}$ from
-$p(\eta|\mathrm{rest})\propto\Psi(u,k)p(\eta),\quad\Psi(u,k):=\sum_{m=0}^{\infty}\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k).$
+$p(\eta\mid\mathrm{rest})\propto\Psi(u,k)p(\eta),\quad\Psi(u,k):=\sum_{m=0}^{\infty}\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k).$
+Step 3’. Sample the hyperparameter $\rho$ of the $h$ from
+$p(\rho\mid\mathrm{rest})\propto\Psi(u,k)\prod_{j=1}^{k}\kappa(u;n_{j})p(\rho).$
 Step 4. Sample $M_{na}$ from the following discrete distribution
-$\mathbb{P}(M_{na}=m|\mathrm{rest})\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=0,1,\dots.$
-[…]
-[…]
-$p(\tau_{m}|\mathrm{rest})\propto\left\{\prod_{i;\theta_{i}=\tau_{m}}f(y_{i}|\tau_{m})\right\}p(\tau).$
-  Step 7. Sample $S_{m_{na}}$, for $\ m_{na}=k+1,\dots,M_{na}+k$from $\mathrm{GIG}(2u+1,\alpha^{2},-1/2).$
-[…]
-[…]
+$\mathbb{P}(M_{na}=m\mid\mathrm{rest})\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=0,1,\dots.$
+Step 5. Sample $S_{m}$, for $m=1,\dots,k$ from $e^{-S_{m}u}S_{m}^{n_{m}}h(S_{m}).$
+Step 6. Sample $\tau_{m}$, for $m=1,\dots,k$ from
+$p(\tau_{m}\mid\mathrm{rest})\propto\left\{\prod_{i;\theta_{i}=\tau_{m}}f(y_{i}\mid\tau_{m})\right\}p(\tau).$
+Step 7. Sample $S_{m_{na}}$, for $m_{na}=k+1,\dots,M_{na}+k$ from $e^{-S_{m}u}h(S_{m}).$
+Step 8. Sample $\tau_{m_{na}}$, for $m_{na}=k+1,\dots,M_{na}+k$ from the prior $p(\tau)$.
+We summarize the algorithm in Algorithm 1.
+The point of this algorithm is the data augmentation through the latent variable such as $U_{n}\mid T\sim\mathrm{Gamma}(n,T)$, where $n$ is the sample size and $T=\sum_{m=1}^{M}S_{m}$.
+It is important to sample the number of empty components $M_{na}$ from (2.7) in Step 4.
+This step allows for direct sampling of $M$ by adding $k$ and label variables to be updated as in the finite mixture model with given $M$ in step 2.
+The update is the same as the telescoping sampling proposed by fruhwirth2021generalized, and the method is more efficient than the classical restaurant process.
+However, in implementation, it is important to determine $q_{M}$ so that the series $\Psi(u,k)$ in Steps 3 and 4 can be written analytically and random variables can be easily generated from the full conditional distributions of $\eta$ and $M_{na}$.
+In Step 5 and 7, the allocated and unallocated weights are updated based on (2.5) and (2.6), respectively.
+The prior $h$ of $S_{m}$ is discussed in detail in the next paragraph.
+In Steps 7, 8, the assigned unnormalized weights and kernel parameters are updated.
+When $M_{na}$ sampled in Step 4 is greater than or equal to $1$, Steps 7 and 8 are executed.
+To implement the blocked Gibbs sampling, at least two conditions must be satisfied: 1) both the probability density function and the Laplace transform must be available in closed form, 2) each $S_{m}$ can be easily generated from (2.5) and (2.6).
+Additionally, it is desirable that the cumulants of $h$ can be expressed in closed form when updating $\eta$ or when extending the (static) MFM to the DMFM.
+When $S_{m}$ follows the $\sigma$-stable distribution, the blocked Gibbs sampling, in general, cannot be used, as the probability density function is available in integral form.
+Alternatively, a marginal Gibbs sampler based on the restaurant process is required, as employed in argiento2022infinity.
+The inverse Gaussian distribution satisfies all of the conditions.
+The probability distribution, the Laplace transform and the cumulant are given by (3.1), (3.3) and (3.4) respectively.
+If $h$ is the $\text{IGau}(\alpha,1)$, the full conditional distribution of $S_{m}$ is the generalized inverse Gaussian distribution.
+From (2.5) and (2.6), we have
+$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{ind}}{\sim}\mathrm{GIG}(2u+1,\alpha^{2},n_{m}-1/2),\quad m\in\mathcal{M}_{a},$
+$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{i.i.d}}{\sim}\mathrm{GIG}(2u+1,\alpha^{2},-1/2),\quad m\in\mathcal{M}_{na},$
+where $\mathrm{GIG}(a,b,c)$ is the generalized inverse Gaussian distribution with $a>0,b>0,c\in\mathbb{R}$.
 3.3 Prior distributions for the number of mixture components
-Assume that $M-1$ follows a discrete probability distribution with the support $\{0,1,2,\dots\}$.
-In this paper, we consider $\text{Poisson}(\Lambda)$ ($\Lambda>0$) for $M-1$, because the constraints in Steps 3 and 4 are satisfied.
-In fact, from Argiento and De Iorio (2022), we have
+Assume that M1M-1 follows a discrete probability distribution with the support $\{0,1,2,\dots\}$.
+In this paper, we consider $\text{Poisson}(\Lambda)$ ($\Lambda>0$) for M1M-1, because the constraints in Steps 3 and 4 are satisfied.
+In fact, from argiento2022infinity, we have
 $\Psi(u,k)=\Lambda^{k-1}(\Lambda\psi(u)+k)\exp(\Lambda(\psi(u)-1)).$ (3.6)
 Assuming $\Lambda\sim\text{Gamma}(a_{\Lambda},b_{\Lambda})$ for $a_{\Lambda},b_{\Lambda}>0$, the full conditional distributions of $\Lambda$ and $M_{na}$ are given by
-$p(\Lambda|\mathrm{rest})\propto\psi(u)(k+a_{\Lambda}-1)$ $\mathrm{Ga}(k+a_{\Lambda}+1,1-\psi(u)+b_{\Lambda})$
-$+k(b_{\Lambda}+1-\psi(u))\mathrm{Ga}(k+a_{\Lambda},1-\psi(u)+b_{\Lambda}),$
-[…]
-[…]
+$p(\Lambda\mid\mathrm{rest})\propto\psi(u)(k+a_{\Lambda}-1)$ $\mathrm{Gamma}(k+a_{\Lambda}+1,1-\psi(u)+b_{\Lambda})$
+$+k(b_{\Lambda}+1-\psi(u))\mathrm{Gamma}(k+a_{\Lambda},1-\psi(u)+b_{\Lambda}),$
+$\mathbb{P}(M_{na}=m\mid\mathrm{rest})\propto\Lambda\psi(u)$ $\textrm{Shifted-Poisson}(\Lambda\psi(u),1)+k\mathrm{Poisson}(\Lambda\psi(u)),$
+respectively, where $\text{Shifted-Poisson}(\Lambda,t)$ is parallel shifted of the distribution $\text{Poisson}(\Lambda)$ by $t$.
+Furthermore, if $b_{\Lambda}$ is hierarchically assigned a beta-prime prior $\text{Betaprime}(a_{p},b_{p}),a_{p},b_{p}>0$, then the marginal prior of M1M-1 follows a beta-negative binomial distribution $\text{BNB}(a_{\Lambda},a_{p},b_{p})$.
+fruhwirth2021generalized recommend using $\text{BNB}(1,4,3)$ under Dirichlet weights.
+It should be noted, however, that updating $b_{\Lambda}$ requires the Metropolis–Hastings algorithm.
+Therefore, various discrete distributions can be incorporated into the prior for M1M-1 by using a Poisson-based hierarchical structure.
 3.4 Specification of kernels
@@
 3.4.1 Cluster analysis and density estimation
-One of the most famous and useful kernels is the (multivariate) normal kernel $N(\cdot|\mu,\sigma^{2})$.
-In the univariate case, we often use the normal inverse gamma model $N(\mu|m_{0},\sigma^{2}/\eta)\times\text{IG}(\sigma^{2}|c_{0},C_{0})$ as a prior distribution of $\xi=(\mu,\sigma^{2})$, where $\ m_{0}\in\mathbb{R}$, $\eta,\ c_{0},\ C_{0}>0$.
+One of the most famous and useful kernels is the (multivariate) normal kernel $N(y\mid\mu,\sigma^{2})$.
+In the univariate case, we often use the normal inverse gamma model as a prior distribution of $\tau=(\mu,\sigma^{2})$: $\mu\mid\sigma^{2}\sim N(m_{0},\sigma^{2}/\eta),\quad\sigma^{2}\sim\text{IG}(c_{0},C_{0})$, where $m_{0}\in\mathbb{R}$, $\eta,\ c_{0},\ C_{0}>0$.
 The parameter $\eta$ is called a smoothing parameter and plays an important role in density estimation.
 It is possible to include a hierarchical prior for $\eta$.
-In the multivariate case, we often employ the normal inverse Wishart model $N_{r}(\mu|m_{0},\Sigma/B_{0})\times\mathrm{W}^{-1}(\Sigma|c_{0},C_{0})$ as a prior distribution of $(\mu,\Sigma)$, where $\ m_{0}\in\mathbb{R}^{r},\ B_{0}>0,\ c_{0}>r-1$ and $C_{0}$ is a positive definite matrix.
-We will use normal kernels in later numerical experiments in Sections 4.1 and 4.2 for clustering and density estimation.
-In the context of finite mixture models, various kernels have been proposed.
-If we have prior information that the data have skewness, the skew normal or skew-t kernel is also useful (see e.g., Frühwirth-Schnatter and Pyne, 2010).
-The skew-normal and skew-t distributions is easy to handle because these kernels have the scale mixtures of normal representations, except for the degree of freedom of skew-t distribution.
+In the multivariate case, the hierarchical normal inverse Wishart model: $\mu\sim N_{r}(b_{0},B_{0}),\Sigma^{-1}\mid C\sim W(c_{0},C),C\sim W(g_{0},G_{0})$, where $b_{0}\in\mathbb{R}^{r}$, $B_{0}$ is an $r\times r$ covariance matrix, $c_{0},g_{0}>r-1$, and $G_{0}$ is an $r\times r$ positive definite symmetric matrix.
+We used normal kernels in later numerical experiments in Sections 4.1 and 4.2 for clustering and density estimation.
+Of course, other kernels can also be used depending on the purpose (see e.g., fruhwirth2010bayesian).
 3.4.2 Network analysis
 As an application of the proposed method, we perform community detection on network data.
-Community detection is the task of identifying dense subclasses in network data and corresponds to clustering and estimating the number of components, called the number of communities in network analysis.
+Community detection is the task of identifying dense subclasses in network data, corresponds to clustering and estimating the number of components, called the number of communities in network analysis.
 Note that the number of components of finite mixture models is equivalent to the number of communities in the network, and both are denoted $M$.
-Estimating the number of communities is an important problem and various methods have been proposed (Shi and Malik, 2000; Newman, 2004; White and Smyth, 2005).
-From a model-based perspective, it is essentially the same as estimating the number of components in a finite-mixture model, and we can apply MFM.
-The stochastic block model is a famous statistical model of network data (Henze, 1986; Nowicki and Snijders, 2001; Geng et al., 2019), which assumes a stochastic block structure behind and specifies the community structure by estimating the probability of edges being drawn between each group.
-We estimate the number of communities with the MFM in the framework of this stochastic block model.
-Geng et al.
-(2019) proposed a stochastic block model based on MFM with Dirichlet weights, and also constructed a similar algorithm to Miller and Harrison (2018).
+Estimating the number of communities is an important problem and various methods have been proposed (shi2000normalized; newman2004detecting; white2005spectral).
+From a model-based perspective, it is essentially the same as estimating the number of components in a finite mixture model, and we can apply MFM.
+The stochastic block model is a famous statistical model of network data (holland1983stochastic), which assumes a stochastic block structure behind and specifies the community structure by estimating the probability of edges being drawn between each group.
+geng2019probabilistic proposed a stochastic block model based on MFM with Dirichlet weights, and also constructed a similar algorithm to miller2018mixture.
 MFM can be easily applied to community detection by modifying the kernel.
-Data $y$ are replaced by the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $A=(A_{ij})\in\{0,1\}^{n\times n}$ and $n$ is the size of the node.
+Data $y$ are replaced by the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $A=(A_{ij})\in\{0,1\}^{n\times n}$ and $n$ is the number of the node.
 When $A_{ij}=1$, this indicates that an edge is drawn from the $i$ th node to the $j$ th node, and when $A_{ij}=0$, this does not.
-For simplicity, we assume that the adjacency matrix is not direct and does not have a self-loop, in that $A_{i,j}=A_{ji}$ and $A_{ii}=0$, where $1\leq i<j\leq n$.
+For simplicity, we assume that the adjacency matrix is not directed and does not have a self-loop, in that $A_{ij}=A_{ji}$ and $A_{ii}=0$, where $1\leq i<j\leq n$.
 The stochastic block model is formulated as follows,
-$\begin{split}&A_{ij}|Q,M\sim\text{Bernoulli}(\theta_{ij}),\quad\theta_{ij}=Q_{c_{i}c_{j}},\quad 1\leq i<j\leq n,\\ &Q_{rs}\sim\text{Beta}(a_{Q},b_{Q}),\quad 1\leq r\leq s\leq M,\end{split}$ (3.5)
+$\begin{split}&A_{ij}\mid Q,M\sim\text{Bernoulli}(\theta_{ij}),\quad\theta_{ij}=Q_{c_{i}c_{j}},\quad 1\leq i<j\leq n,\\ &Q_{rs}\sim\text{Beta}(a_{Q},b_{Q}),\quad 1\leq r\leq s\leq M,\end{split}$ (3.7)
 where $a_{Q},\ b_{Q}>0$ and $M$ is the number of communities.
@@
 Each element $Q_{rs}$ represents the probability that an edge is drawn between any node belonging to the community label $r$ and any node belonging to the community label $s$.
-To perform community detection based on the proposed model, we just set the Bernoulli likelihood as the kernel.
+To perform community detection for the proposed model, we just set the Bernoulli likelihood as the kernel.
 3.5 Evaluation of the number of empty components
-From the point of view of the interpretability of the model, the generalization performance, and the computational efficiency, it is reasonable that the number $M_{na}$ should be small.
-For the full conditional distribution of $M_{na}$, the following inequality holds, where $M-1|\Lambda\sim\mathrm{Poisson}(\Lambda)\ \text{and}\ \Lambda\sim\mathrm{Gamma}(a_{\Lambda},b_{\Lambda})\ \text{for}\ a_{\Lambda},b_{\Lambda}>0$.
+For the full conditional distribution of $M_{na}$, the following inequality holds, where $M-1\mid\Lambda\sim\mathrm{Poisson}(\Lambda)\ \text{and}\ \Lambda\sim\mathrm{Gamma}(a_{\Lambda},b_{\Lambda})\ \text{for}\ a_{\Lambda},b_{\Lambda}>0$.
 Proposition 3.1.
 For the full conditional distribution of $M_{na}$, the inequalities
-$\mathbb{P}(M_{na}\geq 1|U_{n}=u,\ k,\ M=m,\ \Lambda)$ $\leq\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right),$ (3.6)
-$\mathbb{P}(M_{na}\geq 1|U_{n}=u,\ k,\ M=m)$ $\leq\psi(u)\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right)$ (3.7)
-holds, where $\psi$ is the Laplace transform of the probability distribution for $h$.
-Proof.
-When $m\neq 0$,
-$\mathbb{P}(M_{na}=m|U_{n}=u,\ k,\ M=m,\ \Lambda)=\frac{k}{\Lambda\psi(u)+k}\mathcal{P}_{0}(m;\Lambda\psi(u))+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\mathcal{P}_{1}(m;\Lambda\psi(u)).$
-Thus, the conditional expectation of $M_{na}$ is
-$\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m,\ \Lambda]$ $=\frac{k}{\Lambda\psi(u)+k}\times\Lambda\psi(u)+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\times(\Lambda\psi(u)+1)$
-$=\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right).$
-Furthermore,
-$\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m]$ $=\mathbb{E}_{\Lambda}[\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m,\ \Lambda]]$
-$=\int\left(\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right)\right)\frac{b_{\Lambda}^{a_{\Lambda}}}{\Gamma(a_{\Lambda})}\Lambda^{a_{\Lambda}-1}\exp(-b_{\Lambda}\Lambda)\mathrm{d}\Lambda$
-$\leq\psi(u)\times\frac{a_{\Lambda}}{b_{\Lambda}}+\frac{\psi(u)}{k}\times\frac{a_{\Lambda}}{b_{\Lambda}}$
-$=\psi(u)\times\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right).$
-Finally, the result follows from Markov’s inequality. ∎
-[…]
-$\psi_{\text{Inv-Ga}}(u)=\exp\left(\alpha\left(1-\sqrt{1+2u}\right)\right),\quad u\geq 0,$ (3.8)
-$\psi_{\text{Ga}}(u)=\left(\frac{\eta}{\eta+u}\right)^{\gamma},\quad u\geq 0.$ (3.9)
-Laplace transforms $\psi_{\text{Inv-Ga}}$ and $\psi_{\text{Ga}}$ are decreasing functions with respect to $u$.
+$\mathbb{P}(M_{na}\geq 1\mid U_{n}=u,\ k,\ \Lambda)$ $\leq\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right),$ (3.8)
+$\mathbb{P}(M_{na}\geq 1\mid U_{n}=u,\ k)$ $\leq\psi(u)\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right)$ (3.9)
+hold, where $\psi$ is the Laplace transform of the density $h$.
+The proof of the proposition is given in the Supplementary Material.
+Proposition 3.1 shows that $\psi(u)$ plays an important role in generating empty components.
+The critical difference between the inverse Gaussian and gamma distributions is the Laplace transform.
+Let $\psi_{\text{Gamma}}(u)$ be the Laplace transforms of $\text{Gamma}(\gamma,1)$:
+$\psi_{\text{Gamma}}(u)=\left(\frac{1}{1+u}\right)^{\gamma},\quad u\geq 0.$ (3.10)
+Laplace transforms $\psi_{\text{IGau}}$ and $\psi_{\text{Gamma}}$ are decreasing functions with respect to $u$.
 The former has exponential decay, while the latter is polynomial.
 Figure 1 shows the graphs of the Laplace transforms of inverse Gaussian and gamma distributions when the shape parameters are $\alpha,\gamma=1.0,0.2,10^{-1},10^{-2},10^{-3}$ and the scale parameters are $1$.
-It can be seen that $\psi_{\text{Inv-Ga}}$ decreases much faster than $\psi_{\text{Ga}}$.
-The speed of this decrease has a significant impact on the appearance of the empty component in estimating the number of components.
-[…]
+It can be seen that $\psi_{\text{IGau}}$ decreases much faster than $\psi_{\text{Gamma}}$, contributing to the suppression of $M_{na}$.
+Figure 1: When shape parameters are $\alpha=\gamma=1,0.2,10^{-1},10^{-2},10^{-3}$, left graphs are Laplace transforms of $\text{IGau}(\alpha,1)$, right graphs are Laplace transforms of $\text{Gamma}(\gamma,1)$.
 Inequalities (3.8) and (3.9) are not for marginal posterior distributions, but for full conditional distributions of $M_{na}$.
 This is due to the fact that the marginal posterior distributions of $U_{n}$ are analytically intractable.
+3.6 Relationship with the telescoping sampling
+In this section, we clarify the connection between the blocked Gibbs sampling and the telescoping sampling.
+The blocked Gibbs sampling is derived from the posterior distribution of a random measure induced by the Norm-IFPP, whereas the telescoping sampling is a partially marginalized algorithm constructed from the EPPF.
+Although the two methods arise from fundamentally different formulations, they share several similarities, including the sampling of $M$.
+To the best of our knowledge, their relationship has not been clarified so far, as both methods were proposed independently around the same time.
+First, we explain that the blocked Gibbs sampling can be derived in a manner similar to the telescoping sampling scheme.
+Specifically, it can be derived from both the full mixture posterior $p(M,S,U_{n},\tau_{1:M},c\mid y)$ and the collapsed mixture posterior $p(M,\mathcal{C},U_{n},\tau_{1:k}\mid y)$, where $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ is a partition of $\{1,\dots,n\}$ into $k$ blocks, where $\mid\mathcal{C}_{i}\mid=n_{i}$ for $i=1,\dots,k$ and $\sum_{i=1}^{k}n_{i}=n$, and the hyperparameters $\rho$ and $\eta$ are omitted for simplicity.
+This full mixture posterior is used to sample $\tau_{1:M}$, $c$, $S$ and $U_{n}$ (Step 1, 2, 5, 6, 7 in Algorithm 1).
+The collapsed mixture posterior is obtained by marginalizing out $S$, the unallocated parameters, and all assignments $c$ that induce the same partition $\mathcal{C}$ in the full mixture posterior.
+Thus, by Theorem 4.1 in argiento2022infinity, the conditional distribution of $M$ given $U_{n}=u,\mathcal{C}$ is
+$\mathbb{P}(M=m\mid\mathcal{C},U_{n}=u)\propto\frac{m!}{(m-k)!}\psi(u)^{m-k}q_{M}(m),\quad m=k,k+1,\dots.$
+This full conditional distribution of $M$ is equivalent to the equation (2.7), shifted by $k$.
+The latent variable $U_{n}$ is essential distinction between the two methods, while the blocked Gibbs sampling shares many advantages with the telescoping sampling discussed in fruhwirth2021generalized.
+In other words, the blocked Gibbs sampling can be regarded as a telescoping sampling augmented with the latent variable $U_{n}$.
+By this latent variable, the full conditional distribution $M_{na}$ (or $M$) can be easily sampled when $q_{M}$ is Poisson distribution or negative binomial distribution.
+This is a major advantage over the telescoping sampling, because, in practice, an upper bound $M_{\mathrm{max}}$ is required and $M$ is sampled from a multinomial distribution over $\{k,\dots,M_{\mathrm{max}}\}$ due to the complex form of $M$’s full conditional distribution.
+3.7 Extended telescoping sampling
+In this section, we extend the telescoping sampling so that it can be applied to model (2.3), thereby relaxing the Dirichlet restriction in fruhwirth2021generalized.
+The conditional distributions for $\mathcal{C}$, given $M=m$ and $U_{n}=u$, and for $M$, given $\mathcal{C},U_{n}=u$ are required to construct our algorithm under (2.3).
+Proposition 3.2.
+Let $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ be a set partition with $n_{j}=|\mathcal{C}_{j}|$ and $\sum_{j=1}^{k}n_{j}=n$. Under (2.3), the EPPF is the following:
+$p(\mathcal{C})=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\sum_{m=k}^{\infty}\frac{m!}{(m-k)!}\psi(u;m)^{m-k}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}q_{M}(m)\mathrm{d}u,\quad$ (3.11)
+where $\psi(u;m)$ and $\kappa(u;n,m)$ are the Laplace transform and cumulant of $h_{M}$, respectively.
+Moreover, the conditional probability of $M=m$, for $m=k,k+1,\dots,$ is proportional to the following:
+$\mathbb{P}(M=m\mid U_{n}=u,\mathcal{C})\propto\frac{m!}{(m-k)!}\psi(u;m)^{m-k}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}q_{M}(m).$ (3.12)
+From Proposition 3.2, the Algorithm 2 is immediately derived, showing only the parts that differ from Algorithm 1.
+If $h_{M}$ is $\text{Gamma}(\gamma/M,1)$ or $\text{IGau}(\alpha/M,1)$, Algorithm 2 can be applied to the DMFM based on the Dirichlet and normalized inverse Gaussian distributions, respectively.
+The main difference from Algorithm 1 is that the Laplace transform and the cumulant of $h_{M}$ depend on $M$.
+Since the full conditional distributions of $M$ (or $M_{na}$) and the summation $\sum_{m=k}^{\infty}\frac{m!}{(m-k)!}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;n_{j},m)^{m-k}q_{M}(m)$ strongly depend on $\psi(u;m)$ and $\kappa(u;n,m)$, they cannot, in general, be expressed in a simple form.
+Thus, Step 3 in Algorithm 2 is replaced by sampling from the full mixture posterior, and $M$ is not marginalized out in Step 3’, as in fruhwirth2021generalized.
+In addition, in Step 4, it is practically necessary to set an upper bound $M_{\text{max}}$ and sample from a multinomial distribution.
+Algorithm 2 Extended telescoping sampling
+Step 3. Sample the hyperparameter $\eta$ of the $q_{M}$ from
+$p(\eta\mid\mathrm{rest})\propto p(\eta)q_{M}(m).$
+Step 3’. Sample the hyperparameter $\rho$ of the $h$ from
+$p(\rho\mid\mathrm{rest})\propto\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}p(\rho).$
+Step 4. Sample $M$ from the following discrete distribution
+$\mathbb{P}(M=m\mid\mathrm{rest})\propto\frac{m!}{(m-k)!}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}q_{M}(m),\quad m=k,k+1,\dots.$
 4 Empirical demonstrations
-We evaluate the performance of the MFM-Inv-Ga and MFM-Ga methods through some numerical experiments.
-Recall that $\gamma$ is the shape parameter of the prior distribution of unnormalized weights in the proposed method (MFM-Inv-Ga), while $\alpha$ is that of MFM-Ga.
-4.1 Inference for the number of mixture components and clustering
-In this subsection, we illustrate the performance of clustering and inference for the number of components using artificial and real data.
+We evaluate the performance of the MFM-IGau and MFM-Ga methods through some numerical experiments.
+Recall that $\gamma$ is the shape parameter of the prior distribution of unnormalized weights in the proposed method (MFM-IGau), while $\alpha$ is that of MFM-Ga.
+4.1 Clustering performance
+In this subsection, we compare the clustering performance of the proposed methods, MFM-IGau and DMFM-IGau, with MFM-Ga and DMFM-Ga, using artificial and real data.
 4.1.1 Artificial data
-In this simulation, we assume that $M_{\mathrm{true}}=3$. We generate data from the following multivariate normal distribution:
-$f(y|\mu_{1},\mu_{2},\mu_{3})=0.8\cdot N_{2}(y|\mu_{1},I_{2})+0.1\cdot N_{2}(y|\mu_{2},I_{2})+0.1\cdot N_{2}(y|\mu_{3},I_{2}),$
-[…]
-  • The posterior mean of the number of components $M$.
-  • The posterior mean of the rand index. The rand index is a measure of the clustering fitting, and the value takes $[0,1]$. When it is close to $1$, the assignment estimate is reasonable.
-  • The posterior probability that the number of empty components $M_{na}$ is equal to 00.
-The respective averages over $50$ repetitions are denoted by $\widehat{\mathrm{RI}}$ and ^$\widehat{\mathbb{P}}(M_{na}=0)$.
-We report the posterior mean of $M$ in Figure 2.
-It is observed that the results of the MFM-Inv-Ga method remain almost the same even if the shape parameter is varied.
-However, the MFM-Ga method tends to overestimate the number of mixture components as the shape parameter decreases.
-Figure 2: Box plots of the posterior mean of the number of components $M$ for each shape parameter. From left to right, shape parameters are set as $\gamma,\alpha=1,0.2,10^{-1},10^{-2},10^{-3}$.
-Table 1 shows the results of the clustering performance and the posterior probability that $M_{na}$ is equal to 00.
-For clustering, the MFM-Inv-Ga and MFM-Ga methods are comparable.
-Both methods have reasonable clustering accuracy, only slightly better for MFM-Ga than for MFM-Inv-Ga.
-As the shape parameters $\alpha$ and $\gamma$ decrease, $\widehat{\mathrm{RI}}_{i}$ increases slightly and its standard deviation decreases.
-This is natural given that for data that have components with relatively very large cluster sizes, a suitable prior on a simplex is one with large mass at the edges or vertices.
-It is important to note that Figure 2 and Table 1 show that the MFM-Inv-Ga method produces few empty components for all scenarios and provides reasonable estimates of $M$, but not MFM-Ga.
-Table 1: Results of the rand index $\widehat{\mathrm{RI}}_{i}$ and posterior probability that the number of empty components is equal to zero ^$\widehat{\mathbb{P}}(M_{na}=0)$ (their standard deviations are shown in parentheses) averaged over 50 Monte Carlo replications for each shape parameters.
-$\widehat{\mathrm{RI}}_{i}$
-$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
-MFM-Inv-Ga $0.993$ $0.994$ $0.995$ $0.995$ $0.995$
-$(0.013)$ $(0.006)$ $(0.006)$ $(0.004)$ $(0.005)$
-MFM-Ga $0.995$ $0.995$ $0.995$ $0.996$ $0.996$
-$(0.006)$ $(0.004)$ $(0.003)$ $(0.004)$ $(0.003)$
-^$\widehat{\mathbb{P}}(M_{na}=0)$
-$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
-MFM-Inv-Ga $1.000$ $0.995$ $0.982$ $0.947$ $0.952$
-$(0.001)$ $(0.008)$ $(0.032)$ $(0.135)$ $(0.087)$
-MFM-Ga $0.979$ $0.555$ $0.328$ $0.088$ $0.066$
-$(0.008)$ $(0.080)$ $(0.066)$ $(0.012)$ $(0.013)$
-This can be seen in Table 2.
-It can also be seen that MFM-Inv-Ga assigns a very high posterior probability to $M_{\mathrm{true}}$ than MFM-Ga, and the behavior of the posterior distributions for $k$ and $M$ is similar, together with Table 1.
-Table 2: Posterior probabilities and their standard deviations (shown in parentheses) of the number of components $M$ averaged over 50 Monte Carlo replications. The highest value is bolded.
-MFM-Inv-Ga
-$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
-$\alpha=1$ $0.000$ $0.000$ $\bf{0.973}$ $0.027$ $0.000$ $0.000$
-$(0.000)$ $(0.000)$ $(0.093)$ $(0.093)$ $(0.000)$ $(0.000)$
-$\alpha=0.2$ $0.000$ $0.080$ $\bf{0.886}$ $0.033$ $0.001$ $0.000$
-$(0.000)$ $(0.274)$ $(0.268)$ $(0.048)$ $(0.003)$ $(0.000)$
-$\alpha=10^{-1}$ $0.000$ $0.000$ $\bf{0.942}$ $0.051$ $0.006$ $0.001$
-$(0.000)$ $(0.000)$ $(0.084)$ $(0.069)$ $(0.013)$ $(0.005)$
-$\alpha=10^{-2}$ $0.000$ $0.020$ $\bf{0.890}$ $0.063$ $0.015$ $0.012$
-$(0.000)$ $(0.141)$ $(0.199)$ $(0.078)$ $(0.039)$ $(0.048)$
-$\alpha=10^{-3}$ $0.000$ $0.050$ $\bf{0.860}$ $0.068$ $0.015$ $0.007$
-$(0.000)$ $(0.209)$ $(0.225)$ $(0.084)$ $(0.029)$ $(0.020)$
-MFM-Ga
-$\gamma=1$ $0.000$ $0.000$ $\bf{0.948}$ $0.049$ $0.002$ $0.000$
-$(0.000)$ $(0.000)$ $(0.055)$ $(0.049)$ $(0.006)$ $(0.001)$
-$\gamma=0.2$ $0.000$ $0.010$ $\bf{0.508}$ $0.301$ $0.121$ $0.061$
-$(0.000)$ $(0.070)$ $(0.093)$ $(0.037)$ $(0.037)$ $(0.035)$
-$\gamma=10^{-1}$ $0.000$ $0.000$ $\bf{0.307}$ $0.306$ $0.196$ $0.191$
-$(0.000)$ $(0.007)$ $(0.093)$ $(0.037)$ $(0.037)$ $(0.035)$
-$\gamma=10^{-2}$ $0.000$ $0.003$ $0.088$ $0.155$ $\bf{0.176}$ $0.578$
-$(0.000)$ $(0.002)$ $(0.021)$ $(0.016)$ $(0.013)$ $(0.071)$
-$\gamma=10^{-3}$ $0.000$ $0.003$ $0.066$ $0.132$ $\bf{0.163}$ $0.636$
-$(0.000)$ $(0.020)$ $(0.019)$ $(0.013)$ $(0.011)$ $(0.072)$
-Table 3 shows the CPU times of MFM-Inv-Ga and MFM-Ga for $50$ repetitions of MCMC with $2000$ iterations, where $\alpha,\gamma=1.0$ and $10^{-3}$.
-When $\alpha=\gamma=1.0$, MFM-Inv-Ga is faster than MFM-Ga on average, but MFM-Inv-Ga has more variability than MFM-Ga.
-On the other hand, MFM-Ga is very time-consuming in $\gamma=10^{-3}$, because many empty components are created.
-As a result, MFM-Inv-Ga has the same clustering accuracy as MFM-Ga and outperforms MFM-Ga in terms of $M$ estimation and CPU time.
-Table 3: Comparison of CPU times (in seconds) and their standard deviations (shown in parentheses) for artificial data averaged over $50$ replications, $\alpha,\ \gamma=1.0,10^{-3}$
-MFM-Inv-Ga MFM-Ga
-$\alpha=\gamma=1$ $88.545$ $90.898$
-$(5.686)$ $(2.774)$
-$\alpha=\gamma=10^{-3}$ $87.061$ $192.234$
-$(12.286)$ $(8.932)$
+In this simulation, we assume that $M_{\mathrm{true}}=k_{\mathrm{true}}=8$, and data are generated from the following multivariate normal distribution mixture:
+$f(y\mid\mu_{1},\dots,\mu_{8})=\sum_{i=1}^{8}w_{i}N_{2}(y\mid\mu_{i},\sigma^{2}I_{2}),$
+where $w_{i}=i^{-1.75}/\sum_{i=1}^{8}w_{i}$, $\mu_{1}=(2,0)^{\top},\mu_{2}=(2,5)^{\top},\mu_{3}=(6,0)^{\top},\mu_{4}=(6,5)^{\top},\mu_{5}=(10,0)^{\top},\mu_{6}=(10,5)^{\top},\mu_{7}=(14,0)^{\top},\mu_{8}=(14,5)^{\top}$, $\sigma^{2}=0.45$ and $I_{2}$ is the $2\times 2$ identity matrix.
+We set $n=50,100,200,300,400$ and generate $50$ datasets for each $n$.
+This data generating process represents a situation in which many mixture components have small weights, so that not all components necessarily contribute to the observed data.
+When $n$ is small, it is difficult to obtain observations from all components, whereas as $n$ increases, it becomes more likely that observations are obtained from every component.
+MCMCs were run for 30,000 iterations, with the first 20,000 discarded as burn-in.
+We assume that $M-1\mid\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Gamma}(1,0.1)$ and employ a multivariate normal kernel with a normal–inverse-Wishart hierarchical prior in Section 3.4.1, where $b_{0}=\mathrm{median(y)},B_{0}=\mathrm{diag}(R_{1}^{2},\dots,R_{d}^{2})$, $R_{j}$ is the range of data $y_{1:n}$ in $j$th dimension, $c_{0}=2.5+(d-1)/2,g_{0}=0.5+(r-1)/2$ and $G_{0}=100g_{0}/c_{0}\mathrm{diag}(1/R_{1}^{2},\dots,1/R_{d}^{2})$.
+Also, we set shape parameters $\alpha,\gamma=1.0$ for MFM-IGau, DMFM-IGau and MFM-Ga, and we assume a prior F distribution $F(6,3)$ to shape parameter DMFM-Ga as in fruhwirth2021generalized.
+To measure performance, we consider the posterior mean of the Adjusted Rand index (ARI) (hubert1985comparing), which is a measure of the goodness of clustering.
+When it is close to $1$, the assignment estimate is reasonable.
+The respective averages over $50$ repetitions are denoted by $\widehat{\mathrm{ARI}}$.
+Overall, figure 2 shows that $\widehat{\mathrm{ARI}}$ of all methods increase toward $1$ as $n$ becomes larger.
+MFM-IGau consistently attains higher clustering accuracy than MFM-Ga, and DMFM-IGau exhibits the best overall performance.
+Although the difference between DMFM-IGau and DMFM-Ga is small, DMFM-IGau does not incorporate learning for $\alpha$.
+This result indicates that stable posterior inference is achieved due to the less informative nature of the normalized inverse Gaussian distribution.
+Figure 2: Line plots of the posterior mean of ARI averaged over $50$ Monte Carlo replications for MFM-Ga, DMFM-Ga, MFM-IGau and DMFM-IGau.
 4.1.2 Thyroid data
-We apply the proposed method to famous thyroid data.
-The data is available from the R package mclust and is well known as benchmark data for clustering.
+We apply the proposed method MFM-IGau and DMFM-IGau to famous thyroid data.
+The main purpose of this analysis is to infer the number of cluster $k$ and the number of components $M$ and examine the impact of choice of $q_{M}$.
+The thyroid data are available from the R package mclust and is well known as benchmark data for clustering.
 The sample size is $215$ and the dimension of the data is $6$.
 The thyroid disease of each patient is included and classified into three categories: Normal, hypo and hyper.
-The number of diseases is $150$, $30$ and $35$, respectively.
-Using these labels as true labels, the main interest is the accuracy of the clustering and whether the number of components is estimated to be $3$.
-We use the same model and shape parameters as in Section 4.1.1.
-We independently run the $5$ MCMC chain with different initial values.
-The number of iterations for each chain is $20000$ and finally we get the $10000$ samples.
-We compare MFM-Inv-Ga and MFM-Ga through the posterior mean of $M$, the posterior probability $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ and the posterior mean of the rand index (RI).
-Table 4 shows that the result of the real data analysis is similar to that of Section 4.1.1.
-MFM-Inv-Ga provides a reasonable estimate of $M$ for all shape parameters by not producing empty components.
-However, MFM-Ga overestimates $M$.
-Furthermore, MFM-Inv-Ga is slightly more accurate than MFM-Ga.
-Table 4: Posterior means of $M$ ($\hat{M}$), posterior probabilities of $M_{na}=0$ ($\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$) and posterior means of rand index (RI) for thyroid data.
-MFM-Inv-Ga
-$\hat{M}$ $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ RI
-$\alpha=1$ $3.200$ $1.000$ $0.895$
-$\alpha=0.2$ $3.213$ $0.994$ $0.895$
-$\alpha=10^{-1}$ $3.225$ $0.987$ $0.894$
-$\alpha=10^{-2}$ $3.222$ $0.987$ $0.895$
-$\alpha=10^{-3}$ $3.489$ $0.960$ $0.893$
-MFM-Ga
-$\hat{M}$ $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ RI
-$\gamma=1$ $3.670$ $0.962$ $0.893$
-$\gamma=0.2$ $4.944$ $0.434$ $0.890$
-$\gamma=10^{-1}$ $4.945$ $0.394$ $0.893$
-$\gamma=10^{-2}$ $7.425$ $0.067$ $0.893$
-$\gamma=10^{-3}$ $7.626$ $0.046$ $0.895$
+The number of diseases are $150$, $30$ and $35$, respectively.
+Using these labels as true labels, the main interest is whether the number of components is estimated to be $3$.
+We use priors Poisson distribution $\mathrm{Poisson}(1)$, geometric distribution $\mathrm{Geo}(0.1)$ and beta negative binomial distribution $\mathrm{BNB}(1,4,3)$ for $M$, and $F$-distribution $F(6,3)$ for $\alpha$.
+Table 1 shows that the posterior modes for MFM-IGau and DMFM-IGau are all equal to $3$.
+These results are consistent with those obtained from DMFM-Ga.
+In addition, the table indicates that both the first and third quartiles are $3$, and that the posterior probabilities of $M_{na}=0$ are very close to 1.
+This implies that the difference between the posterior number of components and the posterior number of clusters obtained by MFM-IGau and DMFM-IGau is extremely small.
+Furthermore, the interquartile range of $M$ is tighter than that reported in fruhwirth2021generalized, providing stronger evidence that $M=3$ under our proposed methods than under DMFM-Ga.
+In addition, the overall results do not depend on the choice of the prior $q_{M}$, as is the case for DMFM-Ga.
+Table 1: Posterior summaries of $k$ and $M$ for the thyroid data.
+The leftmost values represent the posterior modes, the values in brackets denote the first and third quartiles, and the rightmost values correspond to the posterior probabilities of $M_{na}=0$.
+$p(k\mid\mathrm{data})$
+$\mathrm{Poi}(1)$ $\mathrm{Geo}(0.1)$ $\mathrm{BNB}(1,4,3)$
+MFM-IGau 33 $[3,3]$ 33 $[3,3]$ 33 $[3,3]$
+DMFM-IGau 33 $[3,3]$ 33 $[3,3]$ 33 $[3,3]$
+$p(M\mid\mathrm{data})$
+$\mathrm{Poi}(1)$ $\mathrm{Geo}(0.1)$ $\mathrm{BNB}(1,4,3)$
+MFM-IGau 33 $[3,3]$ $0.997$ 33 $[3,3]$ $0.980$ 33 $[3,3]$ $1.000$
+DMFM-IGau 33 $[3,3]$ $0.993$ 33 $[3,3]$ $0.973$ 33 $[3,3]$ $0.986$
 4.2 Density estimation
-As seen in the previous subsection, the main difference between the MFM-Inv-Ga and MFM-Ga methods is the frequency of occurrence of the empty components.
-The MFM-Ga method can achieve very high clustering accuracy by setting a small $\gamma$ and allowing empty components.
-However, small $\gamma$ not only makes the model difficult to interpret but also degrades the predictive accuracy of the model.
-In this subsection, we examine this phenomenon through density estimation using predictive distributions.
+In this section, we compare MFM-IGau and MFM-Ga using the Galaxy data.
+The aim of this analysis is to examine how sensitive the posterior distribution of $k$ and the density estimation based on the predictive distribution are to the choice of shape parameters.
 We used the famous galaxy dataset, which is a small data set consisting of $82$ velocities (km/sec) of different galaxies.
-The data is widely used in nonparametric Bayesian statistics as a benchmark for density estimation and cluster analysis.
-The details of the data are found in Roeder (1990).
-We set $M-1|\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1/5)$.
-We also employ the univariate normal kernel $N(y|\mu,\sigma^{2})$, and the normal–inverse gamma conjugate prior $N(\mu|m_{0},\sigma^{2}/\tau)\times\text{IG}(\sigma^{2}|c_{0},C_{0})$ as the prior of the parameters in the kernel.
-Moreover, we assume that the prior of $C_{0}$ is $\text{Gamma}(d_{0},D_{0})$, and we set $m_{0}=(\max(\mathrm{data})+\min(\mathrm{data}))/2,\ d_{0}=0.2,\ D_{0}=10/(\max(\mathrm{data})+\min(\mathrm{data}))^{2}$ as Richardson and Green (1997).
+The data are widely used in nonparametric Bayesian statistics as a benchmark for density estimation and cluster analysis.
+The details and a more comprehensive analysis of the data were given in roeder1990density and grun2022many.
+We set $M-1\mid\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1/5)$.
+We also employ the univariate normal kernel $N(y\mid\mu,\sigma^{2})$, and the normal-nverse gamma conjugate prior in Section 3.4.1: $N(\mu\mid m_{0},\sigma^{2}/\tau)\times\text{IG}(\sigma^{2}\mid c_{0},C_{0})$ as the prior of the parameters in the kernel.
+Moreover, we assume that the prior of $C_{0}$ is $\text{Gamma}(d_{0},D_{0})$, and we set $m_{0}=(\max(\mathrm{data})+\min(\mathrm{data}))/2,\ d_{0}=0.2,\ D_{0}=10/(\max(\mathrm{data})+\min(\mathrm{data}))^{2}$ as richardson1997bayesian.
 The parameter $\tau$ controls the smoothness of the estimated density function.
-We assume that the prior of the smoothing parameter $\tau$ is $\text{Gamma}(w,W)$, where $w=0.5$ and $W=50$ as Escobar and West (1995).
+We assume that the prior of the smoothing parameter $\tau$ is $\text{Gamma}(w,W)$, where $w=0.5$ and $W=50$ as escobar1995bayesian.
 Parameters $\tau$ and $C_{0}$ should be carefully learned from the data, since density estimation is sensitive to their choice.
-The MCMC iterations are $100000$ and the first half of the $90000$ samples are not used as a burn-in period.
-The shape parameters are the same as in Section 4.1.
-We evaluated the result of density estimation and posterior probabilities of $k$, $M$, and $M_{na}$ for each shape parameter.
-Figure 3: Results of density estimation for galaxy data using MFM-Inv-Ga (left) and MFM-Ga (right).
+MCMC was run for 100,000 iterations with a burn-in of 90,000.
+Each value of $\alpha=1.0,0.2,10^{-1},10^{-2},10^{-3}$ is considered, and $\gamma$ is set to $\gamma=\alpha$ in each case.
+Note that the first and second moments of $\mathrm{IGau}(\alpha,1)$ and $\mathrm{Gamma}(\gamma,1)$ are both equal to $\alpha$ and $\gamma$, respectively; thus, this corresponds to a moment-matching setting.
+We evaluated the result of density estimation and posterior probabilities of $k$ for each shapes.
 We show the estimated density functions using the posterior means in Figure 3.
-From the figure, it is observed that the shapes of the estimated densities using the MFM-Inv-Ga do not depend on the choice of the shape parameter $\alpha$.
+From the figure, it is observed that the shapes of the estimated densities using the MFM-IGau do not depend on the choice of the shape parameter $\alpha$.
 However, results using the MFM-Ga method seem to be strongly influenced by the choice of the shape parameter $\gamma$.
 For $\gamma=10^{-2}$ and $10^{-3}$, MFM-Ga cannot capture two large peaks in the middle.
-Table 5: Posterior probabilities of the number of components $M$ and clusters $k$ for Galaxy data.
-MFM-Inv-Ga
-$M\leq 3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M=9$ $M\geq 10$
-$\alpha=1.0$ $0.000$ $0.089$ $0.224$ $0.31$ $0.170$ $0.097$ $0.054$ $0.056$
-$\alpha=0.2$ $0.136$ $0.194$ $0.171$ $0.152$ $0.114$ $0.086$ $0.054$ $0.091$
-$\alpha=10^{-1}$ $0.071$ $0.096$ $0.154$ $0.144$ $0.137$ $0.102$ $0.079$ $0.216$
-$\alpha=10^{-2}$ $0.039$ $0.078$ $0.119$ $0.151$ $0.123$ $0.102$ $0.081$ $0.308$
-$\alpha=10^{-3}$ $0.057$ $0.111$ $0.170$ $0.153$ $0.122$ $0.101$ $0.071$ $0.217$
+Figure 3: Results of density estimation for Galaxy data using MFM-IGau (left) and MFM-Ga (right).
+The number of clusters in the galaxy data has been reported as $5$ or $6$ in existing studies.
+From Table 2, in MFM-IGau, the posterior distributions of $k$ have large probabilities at $k=5$ and $6$ for all $\alpha$, and the posterior distributions of $k$ under MFM-IGau are more similar to each other across different value of $\alpha$ than those under MFM-Ga.
+Focusing on the number of clusters $k$, it is interesting that MFM-IGau is more robust than MFM-Ga with respect to the choice of the shape parameter in estimating $k$.
+This suggests that the prior distribution of $k$ based on MFM-IGau is less informative than MFM-Ga, i.e., the same relationship holds for MFM-IGau and MFM-Ga as for normalized inverse Gaussian process and Dirichlet process.
+Table 2: Posterior probabilities of the number of clusters $k$ for Galaxy data.
+MFM-IGau
 $k\leq 3$ $k=4$ $k=5$ $k=6$ $k=7$ $k=8$ $k=9$ $k\geq 10$
@@
 MFM-Ga
-$M\leq 3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M=9$ $M\geq 10$
-$\gamma=1.0$ $0.038$ $0.104$ $0.191$ $0.206$ $0.176$ $0.122$ $0.072$ $0.091$
-$\gamma=0.2$ $0.018$ $0.042$ $0.076$ $0.098$ $0.114$ $0.117$ $0.103$ $0.432$
-$\gamma=10^{-1}$ $0.019$ $0.040$ $0.057$ $0.071$ $0.075$ $0.083$ $0.079$ $0.577$
-$\gamma=10^{-2}$ $0.009$ $0.023$ $0.042$ $0.059$ $0.071$ $0.076$ $0.075$ $0.645$
-$\gamma=10^{-3}$ $0.003$ $0.004$ $0.013$ $0.019$ $0.025$ $0.033$ $0.038$ $0.865$
 $k\leq 3$ $k=4$ $k=5$ $k=6$ $k=7$ $k=8$ $k=9$ $k\geq 10$
@@
 $\gamma=10^{-3}$ $0.983$ $0.018$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
-The number of clusters in the galaxy data has been reported as $5$ or $6$ in existing studies.
-From table 5, in MFM-Inv-Ga, the posterior distributions of $k$ have large probabilities at $k=5$ and $6$ for all $\alpha$ and each posterior distribution induced by MFM-Inv-Ga is more similar than by MFM-Ga.
-For $M$, both MFM-Inv-Ga and MFM-Ga overestimate when $\alpha$ and $\gamma$ are small.
-This can be seen in Table 6.
-The reason is why the data size $n=82$ is small and $M_{na}$ is easily produced.
-Tables 5 and 6 show that the degree of overestimation is much smaller for MFM-Inv-Ga than for MFM-Ga.
-Table 6: Posterior probabilities of the number of empty components is equal to zero for Galaxy data
-$\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$
-$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
-MFM-Inv-Ga $0.854$ $0.654$ $0.480$ $0.347$ $0.443$
-MFM-Ga $0.648$ $0.081$ $0.040$ $0.010$ $0.002$
-Focusing on the number of clusters $k$, it is interesting that MFM-Inv-Ga is more robust for shape parameter estimates $k$ than MFM-Ga.
-This suggests that the prior distribution of $k$ based on MFM-Inv-Ga is less informative than MFM-Ga, i.e., the same relationship holds for MFM-Inv-Ga and MFM-Ga as for normalized inverse Gaussian process and Dirichlet process.
-In Lijoi et al.
-(2005), the prior of $k$ induced by a normalized inverse Gaussian process is not more informative for the precision parameter than the Dirichlet process.
-However, the prior $k$ for the normalized inverse Gaussian process can be written in closed form, while that for MFM-Inv-Ga is given in complex integral form.
 The shape parameter of MFM-Ga should be chosen carefully because it has a significant impact on clustering, density estimation, and the appearance of empty components.
-However, MFM-Inv-Ga is much more robust than MFM-Ga with respect to the choice of the shape parameter.
-Hence, MFM-Inv-Ga is superior to MFM-Ga in that it is much easier to use than MFM-Ga.
+However, MFM-IGau is much more robust than MFM-Ga with respect to the choice of the shape parameter.
+This is consistent with the earlier remark that the normalized inverse Gaussian distribution is less informative and more robust to parameter specification than the Dirichlet distribution, leading to more stable inference.
+Note that results do not differ substantially when information measures such as the expected Gini index are matched.
 4.3 Community detection
@@
 First, we illustrate the performance of the proposed method using simulation data.
-We assume that the true number of communities is $3$, denoted by $M_{\mathrm{true}}=3$ and the number of nodes in the network is set as $n=150$.
-We here consider the balanced network, in that the true allocation consists of $M_{\mathrm{true}}$ communities with $50$ nodes.
-For the true probability matrix $Q$, we assume that each component is expressed by $q_{rs}=q+(p-q)I(r=s)$, where $q=0.1$, $p=0.8$ and $1\leq r\leq s\leq M_{\mathrm{true}}$.
-The assumption indicates that the edges are more easily drawn within the same community and less easily drawn between different communities.
-In the setup, we generate the $50$ data set.
-We set $M-1|\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1)$, and employ (3.5) as a kernel and a prior, where $a_{Q}=1$ and $b_{Q}=1$.
-The values of the shape parameters, the evaluations and the MCMC setting are the same as those of Section 4.1.1.
-Figure 4: Box plots of the posterior mean of the number of communities $M$. From left to right, shape parameters are set as $\gamma,\alpha=1,0.2,10^{-1},10^{-2},10^{-3}$.
-We report the posterior mean of the number of communities in Figure 4. It is observed that the results are almost the same as those of Figure 2.
-Table 7: Results of the rand index $\widehat{\mathrm{RI}}_{i}$ and posterior probability that the number of empty components is equal to zero ^$\widehat{\mathbb{P}}(M_{na}=0)$ (their standard deviations are shown in parentheses) averaged over 50 Monte Carlo replications.
-$\widehat{\mathrm{RI}}_{i}$
-$\alpha=\gamma=1$ $\alpha=\gamma=0.2$ $\alpha=\gamma=10^{-1}$ $\alpha=\gamma=10^{-2}$ $\alpha=\gamma=10^{-3}$
-MFM-Inv-Ga $0.997$ $0.997$ $0.995$ $0.991$ $0.991$
-$(0.010)$ $(0.019)$ $(0.032)$ $(0.044)$ $(0.044)$
-MFM-Ga $0.995$ $0.996$ $0.980$ $0.987$ $0.910$
-$(0.032)$ $(0.026)$ $(0.059)$ $(0.054)$ $(0.044)$
-^$\widehat{\mathbb{P}}(M_{na}=0)$
-$\alpha=\gamma=1$ $\alpha=\gamma=0.2$ $\alpha=\gamma=10^{-1}$ $\alpha=\gamma=10^{-2}$ $\alpha=\gamma=10^{-3}$
-MFM-Inv-Ga $1.000$ $0.994$ $0.988$ $0.975$ $0.962$
-$(0.001)$ $(0.009)$ $(0.029)$ $(0.044)$ $(0.094)$
-MFM-Ga $0.960$ $0.500$ $0.311$ $0.089$ $0.068$
-$(0.011)$ $(0.071)$ $(0.064)$ $(0.024)$ $(0.018)$
-Table 7 also shows that the results of the posterior probability of $M_{na}$ are the same as Table 1.
-However, in terms of clustering accuracy for network data, MFM-Inv-Ga is higher and more accurate than MFM-Ga.
-Table 8: Posterior probabilities and their standard deviations (shown in parentheses) of the number of components $M$ averaged over $50$ Monte Carlo replications. The highest value is bolded.
-MFM-Inv-Ga
-$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
-$\alpha=1$ $0.000$ $0.000$ $\bf{0.927}$ $0.073$ $0.000$ $0.000$
-$(0.000)$ $(0.000)$ $(0.243)$ $(0.243)$ $(0.000)$ $(0.000)$
-$\alpha=0.2$ $0.000$ $0.012$ $\bf{0.966}$ $0.022$ $0.000$ $0.000$
-$(0.000)$ $(0.084)$ $(0.111)$ $(0.073)$ $(0.001)$ $(0.000)$
-$\alpha=10^{-1}$ $0.000$ $0.020$ $\bf{0.950}$ $0.027$ $0.002$ $0.000$
-$(0.000)$ $(0.141)$ $(0.178)$ $(0.106)$ $(0.009)$ $(0.002)$
-$\alpha=10^{-2}$ $0.000$ $0.036$ $\bf{0.936}$ $0.023$ $0.004$ $0.001$
-$(0.000)$ $(0.179)$ $(0.184)$ $(0.035)$ $(0.008)$ $(0.004)$
-$\alpha=10^{-3}$ $0.000$ $0.028$ $\bf{0.938}$ $0.023$ $0.007$ $0.004$
-$(0.000)$ $(0.143)$ $(0.180)$ $(0.044)$ $(0.021)$ $(0.017)$
-MFM-Ga
-$\gamma=1$ $0.000$ $0.020$ $\bf{0.930}$ $0.049$ $0.002$ $0.000$
-$(0.000)$ $(0.139)$ $(0.135)$ $(0.028)$ $(0.002)$ $(0.000)$
-$\gamma=0.2$ $0.000$ $0.010$ $\bf{0.480}$ $0.307$ $0.133$ $0.070$
-$(0.000)$ $(0.069)$ $(0.085)$ $(0.036)$ $(0.035)$ $(0.038)$
-$\gamma=10^{-1}$ $0.000$ $0.035$ $\bf{0.300}$ $0.288$ $0.186$ $0.191$
-$(0.000)$ $(0.104)$ $(0.059)$ $(0.042)$ $(0.040)$ $(0.070)$
-$\gamma=10^{-2}$ $0.000$ $0.010$ $0.091$ $0.156$ $\bf{0.175}$ $0.566$
-$(0.000)$ $(0.041)$ $(0.036)$ $(0.014)$ $(0.015)$ $(0.092)$
-$\gamma=10^{-3}$ $0.000$ $0.006$ $0.069$ $0.131$ $\bf{0.160}$ $0.634$
-$(0.000)$ $(0.028)$ $(0.023)$ $(0.017)$ $(0.012)$ $(0.085)$
-From table 8, MFM-Inv-Ga has a much higher posterior probability at $M_{\text{true}}$ for all shape parameters than MFM-Ga.
-As a result, in the context of community detection, MFM-Inv-Ga also achieves better clustering and community estimation than MFM-Ga.
-We also compared the computation time with Geng’s method (Geng et al., 2019), denoted MFM-Geng, when the shape parameters are $1.0$ and $10^{-3}$.
-The result in Table 9 indicates that MFM-Inv-Ga is on average more than twice as efficient as MFM-Geng in terms of computation time.
-The reason is why our algorithm does not update the label variable based on the restaurant process and MFM-Inv-Ga has a structure that is unlikely to produce empty components.
-The MFM-Geng can estimate the number of clusters $k$, but cannot directly estimate the number of communities $M$.
-Furthermore, it is difficult in the MFM-Geng to estimate the hyperparameter of $q_{M}$, because it is necessary to perform complex series calculations, including its parameter.
-In summary, MFM-Inv-Ga (and MFM-Ga) are superior to MFM-Geng in terms of computation time, direct estimation of $M$, and estimation of an essential parameter.
-Table 9: Comparison of CPU times in seconds (standard deviations are shown in parentheses) averaged over $50$ replications.
-MFM-Inv-Ga MFM-Geng
-$\alpha=\gamma=1$ $38.793$ $87.519$
-$(2.894)$ $(3.719)$
-$\alpha=\gamma=10^{-3}$ $38.361$ $83.423$
-$(4.431)$ $(4.673)$
+We assume that the true number of communities and clusters is $2$, denoted by $M_{\mathrm{true}}=k_{\text{true}}=2$ and the number of nodes in the network is set as $n=50$.
+We generated 50 datasets under this setting.
+We here consider the balanced network, in that the true allocation consists of $M_{\mathrm{true}}$ communities with $25$ nodes.
+For the true probability matrix $Q$, we assume that each component is expressed by $q_{rs}=q+(p-q)I(r=s)$, where $q=0.1$, $p=0.35$ and $1\leq r\leq s\leq M_{\mathrm{true}}$.
+We use MFM-Geng, which is suggested by geng2019probabilistic and equivalent to MFM-Ga, and DMFM-Ga as baseline methods.
+In MFM-Geng, $\gamma$ is fixed at $1$, and $M$ is sampled by the post-preseeding step.
+We assign a $\text{BNB}(1,4,3)$ prior to M1M-1 and an $\mathcal{F}(6,3)$ prior to the shape parameters, employ (3.7) as a kernel and a prior, where $a_{Q}=1$ and $b_{Q}=1$.
+We evaluate the performance in terms of $\widehat{\mathrm{ARI}}$, $\widehat{\mathbb{P}}(M_{na}=0)$ and the following two criteria:
+  • Modularity (Mod) (newman2004finding): This index measures the quality of the estimated community structure in a network.
+We computed this value based on the MAP estimate of the label $c$ obtained by the method of dahl2009modal.
+We reported $\widehat{\mathrm{Mod}}$, averaged over $50$ replications.
+  • MCMC run times (seconds): All computations were implemented in R (version 4.4.0) on a laptop equipped with an Apple M2 chip (8-core CPU) and 16 GB of RAM, running macOS Sonoma 14.5.
+MCMCs were run for 100,000 iterations with a burn-in of 50,000 (no thinning).
+Figure 4: Bar plots of the posterior median number of clusters $k$. From left to right, the models are MFM-IGau, DMFM-IGau, MFM-Geng, DMFM-Ga.
+From Figure 4 and the $\widehat{\mathrm{ARI}}$ values in Table 3, the posterior estimates of $k$ and the clustering accuracy are comparable across all methods.
+Although DMFM-Ga achieves a slightly higher ARI than the other methods, MFM-IGau and DMFM-IGau outperform MFM-Geng and DMFM-Ga in terms of modularity.
+This result indicates that MFM-IGau and DMFM-IGau better capture the community structures given the adjacency matrix.
+From the $\widehat{\mathbb{P}}(M_{na}=0)$ values in Table 3, MFM-IGau and DMFM-IGau support $M=2$ more tightly than the other methods.
+This result is also consistent with Proposition 3.1.
+In Table 3, the most pronounced differences among the methods appear in terms of MCMC run time.
+MFM-IGau has the shortest run time, which can be attributed to the fact that all random numbers required at each step of the MCMC algorithm can be generated straightforwardly.
+DMFM-IGau and DMFM-Ga require more time than MFM-IGau, because sampling $M$ from a multinomial distribution with an upper bound is necessary.
+DMFM-IGau is slower than DMFM-Ga, reflecting the greater complexity of the cumulant of the inverse Gaussian distribution compared with that of the gamma distribution.
+MFM-Geng takes by far the longest time for two reasons: label updates are based on a restaurant–process representation, and, as in DMFM-IGau and DMFM-Ga, $M$ is sampled from a multinomial distribution with an upper bound.
+[…]
+MFM-IGau DMFM-IGau MFM-Geng DMFM-Ga
+$\widehat{\mathrm{ARI}}$ $0.861$ $0.859$ $0.857$ $0.863$
+$(0.090)$ $(0.084)$ $(0.083)$ $(0.087)$
+$\widehat{\mathrm{Mod}}$ $0.256$ $0.259$ $0.249$ $0.254$
+$(0.055)$ $(0.053)$ $(0.068)$ $(0.060)$
+$\widehat{\mathbb{P}}(M_{na}=0)$ $0.988$ $0.976$ $0.938$ $0.908$
+$(0.005)$ $(0.007)$ $(0.013)$ $(0.035)$
+MCMC run time (sec) $192.627$ $301.960$ $626.283$ $236.160$
+$(18.252)$ $(29.480)$ $(33.969)$ $(22.320)$
 4.3.2 Dolphins social network data
-The dolphins social network data is often used as a benchmark.
+The dolphins social network is often used as a benchmark.
 Data can be obtained in http://www-personal.umich.edu/mejn/netdata/.
-The data is constructed as an undirected graph and expresses a small-scale animal social network with $64$ bottlenose dolphins off Doubtful Sound, New Zealand.
+The data is an undirected graph and expresses a small-scale animal social network with $64$ bottlenose dolphins off Doubtful Sound, New Zealand.
 Each node represents a dolphin, and an edge is drawn if two dolphins appear to be closely related to each other.
 In previous studies, it is well-known that the network has two communities.
-The details of the data are found in Lusseau et al.
-(2003).
-As before, we compare the posterior distribution of the number of communities between MFM-Inv-Ga and MFM-Ga, and create a co-clustering matrix of MFM-Inv-Ga as quantifying uncertainty of clustering.
+The details of the data are found in lusseau2003bottlenose.
+We compare the posterior distributions of the numbers of clusters and communities among MFM-IGau, DMFM-IGau, MFM-Geng, and DMFM-Ga.
+In addition, we construct a co-clustering matrix to quantify the uncertainty of clustering and compute the modularity value, which is calculated based on the MAP estimator derived by the method of dahl2009modal.
 In this analysis, we set $M-1\sim\text{Poisson}(1)$ and $a_{Q}=b_{Q}=3.0$.
-Table 10: Posterior probabilities of the number of communities $M$ for dolphins social network data
-MFM-Inv-Ga
-$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
-$\alpha=1$ $0.000$ $0.997$ $0.003$ $0.000$ $0.000$ $0.000$
-$\alpha=0.2$ $0.000$ $0.972$ $0.025$ $0.002$ $0.001$ $0.000$
-$\alpha=10^{-1}$ $0.000$ $0.961$ $0.034$ $0.004$ $0.000$ $0.000$
-$\alpha=10^{-2}$ $0.000$ $0.941$ $0.048$ $0.010$ $0.001$ $0.000$
-$\alpha=10^{-3}$ $0.000$ $0.854$ $0.136$ $0.009$ $0.002$ $0.000$
-MFM-Ga
-$\gamma=1$ $0.000$ $0.949$ $0.049$ $0.002$ $0.000$ $0.000$
-$\gamma=0.2$ $0.000$ $0.605$ $0.307$ $0.074$ $0.012$ $0.001$
-$\gamma=10^{-1}$ $0.000$ $0.514$ $0.340$ $0.117$ $0.025$ $0.004$
-$\gamma=10^{-2}$ $0.000$ $0.619$ $0.303$ $0.069$ $0.000$ $0.000$
-$\gamma=10^{-3}$ $0.370$ $0.400$ $0.172$ $0.048$ $0.007$ $0.001$
-From Table 10, the MFM-Inv-Ga is able to successfully estimate the number of communities $M=2$ regardless of the values of the shape parameter, while the MFM-Ga is not.
-We report a co-cluster matrix of MFM-Inv-Ga for $\alpha=1.0$ and MFM-Geng for $\gamma=1.0$, and the results are shown in Figure 5.
-The result of MFM-Inv-Ga is almost identical to the results reported in Geng et al.
-(2019).
-Furthermore, we confirmed that changing the value of $\alpha$ does not change the co-cluster matrices and the clustering solution based on MAP estimation.
-This shows the efficiency of the proposed method.
-Figure 5: Co-clustering matrix with the MFM-Inv-Ga with $\alpha=1.0$ (left) and the MFM-Geng with $\gamma=1.0$ (right) for dolphins social network data.
+Table 4: Posterior summaries of $k$, $M$, and modularity for the dolphin data.
+The leftmost values represent the posterior modes, the values in brackets denote the 5% and 95% quantiles, and the rightmost values correspond to the posterior probabilities of $M_{na}=0$.
+MFM-IGau DMFM-IGau MFM-Geng DMFM-Ga
+$p(k\mid\mathrm{data})$ 22 $[2,2]$ 22 $[2,2]$ 22 $[2,2]$ 22 $[2,2]$
+$p(M\mid\mathrm{data})$ 22 $[2,2]$ $0.988$ 22 $[2,2]$ $0.974$ 22 $[2,3]$ $0.954$ 22 $[2,3]$ $0.905$
+modularity $0.376$ $0.377$ $0.371$ $0.366$
+From Table 4, although all the results for the number of clusters $k$ are the same, the posterior distributions of $M$ obtained from MFM-IGau and DMFM-IGau support $M=2$ more strongly and more tightly than those obtained from MFM-Geng and DMFM-Ga.
+In addition, from the viewpoint of modularity, MFM-IGau and DMFM-IGau perform slightly better than MFM-Geng and DMFM-Ga.
+Finally, the co-clustering matrices obtained from MFM-IGau and DMFM-IGau are very similar to those from MFM-Geng, indicating that these methods produce reasonable clustering results.
+Figure 5: Co-clustering matrix with the MFM-IGau with $\alpha=1.0$ (left) and the MFM-Geng with $\gamma=1.0$ (right) for dolphins social network data.
 5 Concluding remarks
-We proposed a mixture of finite mixtures (MFM) model based on the normalized inverse Gaussian distribution and constructed an efficient posterior sampling algorithm based on Argiento and De Iorio (2022).
-The proposed method is a finite analog of the inverse Gaussian processes proposed by Lijoi et al.
-(2005).
-We illustrate the performance of the proposed method for clustering, density estimation, and community detection, compared to existing MFM models based on Dirichlet distribution (e.g., Miller and Harrison, 2018; Geng et al., 2019; Argiento and De Iorio, 2022).
-The proposed method is robust against the choice of hyper-parameter $\alpha$, and provided reasonable estimates of the number of components and communities compared to the MFM based on the Dirichlet prior distribution.
-Moreover, the proposed method also has a reasonable predictive performance in the sense of density estimation by suppressing the appearance of empty components.
+We proposed MFM and DMFM based on the normalized inverse Gaussian distribution and constructed an efficient posterior sampling algorithm based on argiento2022infinity and fruhwirth2021generalized.
+We illustrate the performance of the proposed method for clustering, density estimation, and community detection, compared to existing MFM and DMFM models based on the Dirichlet distribution.
 The drawbacks of the proposed method are as follows.
 Some parameters involved in the model do not have closed-form marginal distributions because it is not easy to marginalize out $U_{n}$.
-For example, when we focus on clustering, obtaining the interpretable prior distribution for the number of clusters is very important (see e.g., Zito et al., 2023) to incorporate subjective prior information.
+For example, when we focus on clustering, obtaining the interpretable prior distribution for the number of clusters is very important (see e.g., zito2024bayesian) to incorporate subjective prior information.
 However, the proposed model cannot lead to a tractable marginal prior distribution for the number of clusters.
-In the mixture of finite mixtures, the model consists of a distribution over a simplex based on the normalization of independent random variables.
+Also, MFM models consists of a distribution over a simplex based on the normalization of independent random variables.
 Therefore, the correlation between categories cannot be properly modeled.
@@
 It may be inappropriate for data with positive correlations between categories, such as the proportion of symbiotic organisms present, disease complication data, or gene expression data.
-To address such a problem, it may be necessary to construct the MFM in a more general framework that removes the assumption of independence in the Norm-IFPP by Argiento and De Iorio (2022).
-The construction of MFM models using a more flexible distribution over a simplex that can also express positive correlations such as the generalized Dirichlet distribution (Wong, 1998) is an interesting future topic.
-Furthermore, the proposed model can be applied to spatial data (e.g., Geng et al., 2021) and functional data (e.g., Hu et al., 2023).
-For network data, it is expected to extend the MFM to network with weighted edges, degree-corrected stochastic block models, and mixed membership stochastic block models.
+To address such a problem, it may be necessary to construct the MFM in a more general framework that removes the assumption of independence in the Norm-IFPP by argiento2022infinity.
+The construction of MFMs using a more flexible distribution over a simplex that can also express positive correlations is an interesting future topic.
+Furthermore, the proposed model can be applied to spatial data (e.g., geng2021bayesian), functional data (e.g., hu2023bayesian) and network data with weighted edges, degree-corrected stochastic block models.
 Acknowledgement
-This work was supported by Japan Society for the Promotion of Science, the establishment of university fellowships towards the creation of science technology innovation Grant Number JPMJFS2129.
-This work is partially supported by the Japan Society for the Promotion of Science (grant number: 21K13835).
+The authors would like to thank the editor, associate editor, and two anonymous referees for their careful reading and comments to improve the quality of this article.
+We also thank Naoki Awaya for his valuable comments on an early revision of the manuscript.
+This work was supported by JST SPRING, Grant Number JPMJSP2132, and partially supported by the Japan Society for the Promotion of Science (grant number: 21K13835).
+Appendix
+Appendix A Proofs
+A.1 Proof of the cumulant of $\text{IGau}(\alpha,1)$
+Proof.
+We show by mathematical induction that the following equation holds.
+$(-1)^{n}\frac{d^{n}}{du^{n}}\psi(u)=\alpha^{n}\psi(u)\sum_{s=0}^{n-1}\frac{(n-1+s)!}{s!(n-1-s)!}(1+2u)^{-n/2-s/2}(2\alpha)^{-s},$ (A.1)
+where $\psi(u)=\exp\left(\alpha\left(1-\sqrt{1+2u}\right)\right),u\geq 0$.
+If $n=1$, the left side is $\alpha\psi(u)(1+2u)^{-1/2}$ and is equal to the right side.
+When $n=k$, we assume that (A.1) holds.
+We differentiate both sides of (A.1) by u and $k+1$th derivative is
+$\alpha^{k+1}\psi(u)(1+2u)^{-k/2-1/2}\left(\sum_{s=0}^{k-1}A_{s}(1+2u)^{-s/2}(2\alpha)^{-s}+\sum_{s=0}^{k-1}B_{s}(1+2u)^{-s/2-1/2}(2\alpha)^{-s-1}\right),$
+where
+$A_{s}=\frac{(k-1+s)!}{s!(k-1-s)!},\quad B_{s}=\frac{2(k+s)!}{s!(k-1-s)!},\quad s=0,1,\dots,k-1.$
+$B_{k-1}=(2k)!/k!$ and
+$\sum_{s=0}^{k-1}B_{s}(1+2u)^{-s/2-1/2}(2\alpha)^{-s-1}$ $=\sum_{s=0}^{k-2}B_{s}(1+2u)^{-s/2-1/2}(2\alpha)^{-s-1}+\frac{(2k)!}{k!}(1+2u)^{-k/2}(2\alpha)^{-k}$
+$=\sum_{s=1}^{k-1}B_{s-1}(1+2u)^{-s/2}(2\alpha)^{-s}+\frac{(2k)!}{k!}(1+2u)^{-k/2}(2\alpha)^{-k}.$
+For $s=1,\dots,k-1$,
+$A_{s}+B_{s-1}=\frac{(k+s)!}{s!(k-s)!}$
+holds. Thus,
+$\sum_{s=0}^{k-1}A_{s}(1+2u)^{-s/2}(2\alpha)^{-s}+\sum_{s=0}^{k-1}B_{s}(1+2u)^{-s/2-1/2}(2\alpha)^{-s-1}$
+$=1+\sum_{s=1}^{k-1}(A_{s}+B_{s-1})(1+2u)^{-s/2}(2\alpha)^{-s}+\frac{(2k)!}{k!}(1+2u)^{-k/2}(2\alpha)^{-k}$
+$=\sum_{s=0}^{k}\frac{(k+s)!}{s!(k-s)!}(1+2u)^{-s/2}(2\alpha)^{-s}.$
+(A.1) are derived by this equation, when $n=k+1$. Let $K_{\nu}(z)$ be a modified Bessel function of the second kind. In general,
+$K_{n+1/2}(z)=K_{1/2}(z)\times\sum_{s=0}^{n}\frac{(n+s)!}{s!(n-s)!}(2z)^{-s}.$
+Therefore, (A.1) holds. ∎
+A.2 Proof of Proposition 3.1
+Proof.
+When $m\neq 0$,
+$\mathbb{P}(M_{na}=m\mid U_{n}=u,\ k,\ \Lambda)=\frac{k}{\Lambda\psi(u)+k}\mathcal{P}_{0}(m;\Lambda\psi(u))+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\mathcal{P}_{1}(m;\Lambda\psi(u)).$
+Thus, the conditional expectation of $M_{na}$ is
+$\mathbb{E}[M_{na}\mid U_{n}=u,\ k,\ \Lambda]$ $=\frac{k}{\Lambda\psi(u)+k}\times\Lambda\psi(u)+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\times(\Lambda\psi(u)+1)$
+$=\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right).$
+Furthermore,
+$\mathbb{E}[M_{na}\mid U_{n}=u,\ k]$ $=\mathbb{E}_{\Lambda}[\mathbb{E}[M_{na}\mid U_{n}=u,\ k,\ M=m,\ \Lambda]]$
+$=\int\left(\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right)\right)\frac{b_{\Lambda}^{a_{\Lambda}}}{\Gamma(a_{\Lambda})}\Lambda^{a_{\Lambda}-1}\exp(-b_{\Lambda}\Lambda)\mathrm{d}\Lambda$
+$\leq\psi(u)\times\frac{a_{\Lambda}}{b_{\Lambda}}+\frac{\psi(u)}{k}\times\frac{a_{\Lambda}}{b_{\Lambda}}$
+$=\psi(u)\times\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right).$
+Finally, the result follows from Markov’s inequality. ∎
+A.3 Equivalence of Blocked Gibbs Sampler and Telescoping Sampler
+In this section, we show that the blocked Gibbs sampler is essentially equivalent to the telescoping sampler.
+Our strategy is to derive the computational algorithm under the framework of $(2.1)$ or $(2.4)$ in the same manner as the telescoping sampler.
+To this end, we require the EPPF induced by the Norm-IFPP, which is given below.
+Theorem A.1 (Theorem 4.1 in argiento2022infinity).
+Let $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ be a partition of $\{1,\dots,n\}$ into $k$ blocks, where $|\mathcal{C}_{i}|=n_{i}$ for $i=1,\dots,k$ and $\sum_{i=1}^{k}n_{i}=n$.
+The EPPF $p(\mathcal{C})$ induced under $(2.1)$ or $(2.4)$ is
+$p(\mathcal{C})$ $=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\Psi(u;k)\left\{\prod_{j=1}^{k}\kappa(u;n_{j})\right\}\mathrm{d}u,$
+where,
+$\Psi(u;k)=\sum_{m=0}^{\infty}\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k).$
+Note that, for simplicity, we omit the hyperparameter $\eta$ of $q_{M}$ and the hyperparameter $\rho$ of $h$. The full mixture posterior $p(M,S,\tau_{1:M},c\mid y)$ is
+$p(M,S,\tau,c\mid y)\propto\frac{1}{T^{n}}\left\{\prod_{m\in\mathcal{M}_{a}}S_{m}^{n_{m}}h(S_{m})f(y^{[n_{m}]}\mid\tau_{m})p_{0}(\tau_{m})\right\}\left\{\prod_{m\in\mathcal{M}_{na}}h(S_{m})p_{0}(\tau_{m})\right\}q_{M}(M),$
+where $y^{[n_{m}]}$ denotes the $n_{m}$ observations belonging to the $m$th cluster $\mathcal{C}_{m}$. Also,
+$U_{n}\mid T\sim\mathrm{Gamma}(n,T),\quad\frac{\Gamma(n)}{T^{n}}=\int_{0}^{\infty}e^{-Tu}u^{n-1}\mathrm{d}u.$
+Using the data augmentation with $U_{n}$, we obtain the following:
+$p(M,S,U_{n},\tau_{1:M},c\mid y)$ $\propto\frac{U_{n}^{n-1}}{\Gamma(n)}\left\{\prod_{m\in\mathcal{M}_{a}}f(y_{j}^{[n_{m}]}\mid\tau_{m})p_{0}(\tau_{m})\right\}\left\{\prod_{m\in\mathcal{M}_{a}}S_{m}^{n_{m}}\exp(-S_{m}U_{n})h(S_{m})\right\}$
+$\times\left\{\prod_{m\in\mathcal{M}_{na}}\exp(-S_{m}U_{n})h(S_{m})\right\}q_{M}(M).$
+Marginalizing out the empty components, $S_{1},\dots,S_{M}$, and all allocations $c$ that induce the same partition $\mathcal{C}$ yields the following expression for $p(M,\mathcal{C},U_{n},\tau_{1:k}\mid y)$:
+$p(M,\mathcal{C},U_{n},\tau_{1:k}\mid y)\propto\left\{\prod_{m\in\mathcal{M}_{a}}f(y_{j}^{[n_{m}]}\mid\tau_{m})p_{0}(\tau_{m})\right\}p(\mathcal{C},U_{n}\mid M)q_{M}(M).$
+From Theorem A.1,
+$p(M,\mathcal{C},U_{n},\tau_{1:k}\mid y)\propto U_{n}^{n-1}\left\{\prod_{m\in\mathcal{M}_{a}}f(y_{j}^{[n_{m}]}\mid\tau_{m})p_{0}(\tau_{m})\right\}\frac{M!}{(M-k)!}\left\{\prod_{j=1}^{k}\kappa(n_{j};U_{n})\right\}\psi(U_{n})^{M-k}q_{M}(M).$
+Therefore, we obtain
+$\mathbb{P}(M=m\mid\mathcal{C},U_{n}=u)\propto\frac{m!}{(m-k)!}\psi(u)^{m-k}q_{M}(m),\quad m=k,k+1,\dots.$
+The extended telescoping sampler is immediately obtained in the same manner by applying Proposition 3.2.
+A.4 Proof of Proposition 3.2
+Proof.
+This proposition follows from a slight modification of the proof of Theorem A.1, with most steps carrying over directly.
+Let $\mathcal{A}(\mathcal{C})$ be the set of all labels $c$ that induces partition $\mathcal{C}$, and $c_{1}^{\ast},\dots,c_{k}^{\ast}$ be unique values of $c$.
+The EPPF $p(\mathcal{C})$ is written by $p(\mathcal{C})=\sum_{m=1}^{\infty}p(\mathcal{C}\mid M=m)q_{M}(m)$, since $q_{M}(0)=0$.
+Moreover,
+$p(\mathcal{C}\mid M=m)$ $=\mathbbm{1}_{\{k\leq m\}}\sum_{c\in\mathcal{A}(\mathcal{C})}p(c\mid M=m)$
+$=\mathbbm{1}_{\{k\leq m\}}\sum_{c\in\mathcal{A}(\mathcal{C})}\int p(c\mid\pi,M=m)p(\pi\mid M=m)\mathrm{d}\pi$
+$=\mathbbm{1}_{\{k\leq m\}}\sum_{c\in\mathcal{A}(\mathcal{C})}\mathbb{E}\left[\prod_{j=1}^{k}\pi_{c_{j}^{\ast}}^{n_{j}}\right].$
+Note that $\pi_{c_{j}^{\ast}}^{n_{j}}=(S_{c_{j}^{\ast}}/T)^{n_{j}}$ and $\Gamma(n)=T^{n}=\int_{0}^{\infty}e^{-Tu}u^{n-1}\mathrm{d}u$. Thus,
+$\mathbb{E}\left[\prod_{j=1}^{k}\pi_{c_{j}^{\ast}}^{n_{j}}\right]$ $=\mathbb{E}\left[\frac{1}{T^{n}}\prod_{j=1}^{k}S_{c_{j}^{\ast}}^{n_{j}}\right]$
+$=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\left\{\prod_{j=1}^{k}\int_{0}^{\infty}S_{c_{j}^{\ast}}^{n_{j}}e^{-S_{c_{j}^{\ast}}u}h_{M}(S_{c_{j}^{\ast}})\mathrm{d}S_{c_{j}^{\ast}}\right\}\left\{\prod_{t\notin(c_{1}^{\ast},\dots,c_{k}^{\ast})}\int_{0}^{\infty}e^{-S_{t}u}h_{M}(S_{t})\mathrm{d}S_{t}\right\}\mathrm{d}u$
+$=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\left\{\prod_{j=1}^{k}\mathbb{E}\left[S_{c_{j}^{\ast}}^{n_{j}}e^{-S_{c_{j}^{\ast}}u}\right]\right\}\left\{\prod_{t\notin(c_{1}^{\ast},\dots,c_{k}^{\ast})}\mathbb{E}\left[e^{-S_{t}u}\right]\right\}\mathrm{d}u$
+$=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}\mathrm{d}u,$
+where, if $M=m$, $\psi(u;m)=\int_{0}^{\infty}e^{-uS}h_{M}(S)\mathrm{d}S$ and $\kappa(u;n,m):=(-1)^{n}\frac{d^{n}}{du^{n}}\psi(u;m)=\int_{0}^{\infty}S^{n}e^{-Su}h_{M}(S)\mathrm{d}u$.
+Here, if $M=m$, we have
+$\mathcal{A}(\mathcal{C})=\frac{m!}{(m-k)!k!}\times k!=\frac{k!}{(m-k)!}.$
+Therefore,
+$p(\mathcal{C})$ $=\sum_{m=1}^{\infty}p(\mathcal{C}\mid M=m)q_{M}(m)$
+$=\sum_{m=1}^{\infty}\mathbbm{1}_{\{k\leq m\}}\frac{m!}{(m-k)!}\int_{0}^{\infty}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}\mathrm{d}u$
+$=\sum_{m=k}^{\infty}\frac{m!}{(m-k)!}\int_{0}^{\infty}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}\mathrm{d}u$
+$=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\Omega(u;k)\mathrm{d}u,$
+where,
+$\Omega(u;k):=\sum_{m=k}^{\infty}\frac{m!}{(m-k)!}\psi(u;m)^{m-k}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}q_{M}(m+k).$
+Finally, the conditional distribution of $M$ is derived from the collapsed mixture posterior as follows:
+$p(M,\mathcal{C},u,\tau_{1:k}\mid y)$ $\propto\left\{\prod_{m\in\mathcal{M}_{a}}f(y_{j}^{[n_{m}]}\mid\tau_{m})p_{0}(\tau_{m})\right\}p(\mathcal{C},u\mid M)q_{M}(M)$
+$\propto\left\{\prod_{m\in\mathcal{M}_{a}}f(y_{j}^{[n_{m}]}\mid\tau_{m})p_{0}(\tau_{m})\right\}\frac{m!}{(m-k)!}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}q_{M}(M)$
+$\mathbb{P}(M=m\mid U_{n}=u,\mathcal{C})$ $\propto\frac{m!}{(m-k)!}\psi(u;m)^{m-k}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}q_{M}(m).$
+Appendix B Additional simulation results
+B.1 Gini Index Comparison
+In this section, we compare the Dirichlet distribution and the normalized inverse Gaussian distribution in terms of Gini index.
+Let $x$ be a $d$-dimensional vector on $\mathbb{S}_{d}$. We define the Gini index of $x$ as
+$G(x)=\sum_{i=1}^{d}x_{i}(1-x_{i}).$
+This measure indicates that when $G(x)$ is small, the point $x$ lies near a vertex or an edge of the simplex, whereas when $G(x)$ is large, the point $x$ lies near the barycenter of the simplex.
+We examine the expected values of the Gini index for (symmetry) Dirichlet and normalized inverse Gaussian distribution.
+Let $\pi=(\pi_{1},\dots,\pi_{d})\sim\text{NIGau}(\alpha,\dots,\alpha)$.
+From Proporition 2 in lijoi2005hierarchical,
+$\mathbb{E}[\pi_{i}]=\frac{1}{d},\quad\text{Var}[\pi_{i}]=\alpha^{2}(d-1)e^{d\alpha}\Gamma(-2,d\alpha),\quad i=1,\dots,d,$
+where $\Gamma(\cdot,\cdot)$ is the incomplete gamma function. Thus, we obtain immediately,
+$\mathbb{E}[G(\pi)]=1-\frac{1}{d}-d\alpha^{2}(\alpha-1)e^{d\alpha}\Gamma(-2,d\alpha).$
+From Figure 6, values of the Dirichlet distribution decreases sharply as the parameter approaches the origin.
+In contrast, that of the normalized inverse Gaussian distribution decline much more gradually.
+This corresponds to the fact that, for the Dirichlet distribution with a small $\gamma$, the probability mass tends to concentrate on the vertices or edges of the simplex.
+In contrast, even when $\alpha$ is small, the normalized inverse Gaussian distribution does not induce the same sparsity as the Dirichlet distribution, resulting in a more dispersed distribution over the simplex.
+Therefore, in settings such as the DMFM where the shape parameters naturally become small, using the normalized inverse Gaussian distribution can lead to more stable posterior inference.
+Figure 6: $\mathbb{E}[G(\pi)]$ for $\mathrm{Dirichlet}(\gamma,\gamma,\gamma)$ and $\mathrm{NIGau}(\alpha,\alpha,\alpha)$.
+These values were computed on a grid of 10000 points between 0 and 22 (excluding 0).
+B.2 Additional Numerical Experiment
+We conduct numerical experiments for MFM-IGau and MFM-Ga by matching moments and $\mathbb{E}[G(\pi)]$, focusing on the inference of the number of components.
+We assume that $M_{\mathrm{true}}=k_{\mathrm{true}}=3$, and data are generated from the following:
+$f(y\mid\mu_{1},\mu_{2},\mu_{3})=\sum_{i=1}^{3}w_{i}N_{2}(y\mid\mu_{i},\sigma^{2}I_{2}),$
+where $w_{i}=1/3$, $\mu_{1}=(2,0)^{\top},\mu_{2}=(2,5)^{\top},\mu_{3}=(6,0)^{\top}$, $\sigma^{2}=0.45$ and $I_{2}$ is the $2\times 2$ identity matrix.
+We set $n=300$ and generate $50$ datasets.
+We set $\alpha=0.1,0.01,0.001$ and consider the following two settings for $\gamma$:
+  • Moment match: We set $\gamma=0.1,0.01,0.001$. Note that first and second moments of $\text{Gamma}(\gamma)$ and $\text{IGau}(\alpha,1)$ are $\gamma$ and $\alpha$, respectively.
+  • $\mathbb{E}[G(\pi)]$ match: We set $\gamma=0.490,0.352,0.335$.
+For each $\alpha$, we compute $\mathbb{E}[G(\pi)]$ for $\text{NIGau}(\alpha,\alpha,\alpha)$ and determine the corresponding value of $\gamma$ for each $\alpha$, using the uniroot function in R.
+The remaining settings for kernels are the same as in Section 4.1.1.
+Each MCMC iteration is $30000$ and the first $20000$ samples are not used as a burn-in period.
+We assume that $M-1\mid\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Gamma}(1,1)$.
+In Table 5, we reported posterior probabilities averaged over $50$ repetitions.
+We observe that MFM-IGau consistently yields the highest (averaged) posterior probability at $M=3$.
+On the other hand, in the moment match case, MFM-Ga tends to overestimate $M$, because it produces many empty components when $\gamma$ is small.
+In the Gini index match case, MFM-Ga also assigns the highest posterior probability to $M=3$, similarly to MFM-IGau, but the probability is lower than that of MFM-IGau.
+Moreover, the results for very small $\gamma$, such as $\gamma=0.1,0.01,0.001$, could also be interpreted from the perspective of the Gini index.
+When $\gamma$ is extremely small, the expected Gini index tends to decrease, and consequently the Dirichlet prior assigns very small probabilities to many components.
+Consequently, this prior structure may generate unnecessary empty components.
+Table 5: Posterior probabilities and their standard deviations (shown in parentheses) of the number of components $M$ averaged over $50$ Monte Carlo replications. The highest value is bolded.
+MFM-IGau
+$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M\geq 9$
+$\alpha=10^{-1}$ $0.000$ $0.000$ $\bm{0.971}$ $0.024$ $0.004$ $0.001$ $0.000$ $0.000$ $0.000$
+$(0.000)$ $(0.000)$ $(0.059)$ $(0.046)$ $(0.021)$ $(0.003)$ $(0.000)$ $(0.000)$ $(0.000)$
+$\alpha=10^{-2}$ $0.000$ $0.000$ $\bm{0.959}$ $0.036$ $0.004$ $0.001$ $0.000$ $0.000$ $0.000$
+$(0.000)$ $(0.000)$ $(0.063)$ $(0.054)$ $(0.008)$ $(0.004)$ $(0.002)$ $(0.001)$ $(0.000)$
+$\alpha=10^{-3}$ $0.000$ $0.000$ $\bm{0.941}$ $0.048$ $0.007$ $0.002$ $0.001$ $0.000$ $0.001$
+$(0.000)$ $(0.000)$ $(0.111)$ $(0.096)$ $(0.014)$ $(0.006)$ $(0.003)$ $(0.002)$ $(0.003)$
+MFM-Ga
+$\gamma=10^{-1}$ $0.000$ $0.000$ $\bm{0.331}$ $0.301$ $0.187$ $0.098$ $0.047$ $0.021$ $0.016$
+$(0.000)$ $(0.000)$ $(0.100)$ $(0.021)$ $(0.035)$ $(0.028)$ $(0.017)$ $(0.009)$ $(0.009)$
+$\gamma=10^{-2}$ $0.000$ $0.000$ $0.098$ $0.172$ $\bm{0.187}$ $0.165$ $0.128$ $0.091$ $0.159$
+$(0.000)$ $(0.000)$ $(0.021)$ $(0.024)$ $(0.013)$ $(0.005)$ $(0.009)$ $(0.012)$ $(0.037)$
+$\gamma=10^{-3}$ $0.000$ $0.000$ $0.066$ $0.131$ $0.161$ $0.159$ $0.136$ $0.107$ $\bm{0.240}$
+$(0.000)$ $(0.000)$ $(0.003)$ $(0.004)$ $(0.004)$ $(0.004)$ $(0.003)$ $(0.003)$ $(0.009)$
+$\gamma=0.490$ $0.000$ $0.000$ $\bm{0.798}$ $0.170$ $0.028$ $0.004$ $0.001$ $0.000$ $0.000$
+$(0.000)$ $(0.000)$ $(0.070)$ $(0.048)$ $(0.019)$ $(0.005)$ $(0.001)$ $(0.000)$ $(0.000)$
+$\gamma=0.352$ $0.000$ $0.000$ $\bm{0.717}$ $0.221$ $0.050$ $0.010$ $0.002$ $0.000$ $0.000$
+$(0.000)$ $(0.000)$ $(0.040)$ $(0.024)$ $(0.012)$ $(0.004)$ $(0.001)$ $(0.000)$ $(0.000)$
+$\gamma=0.335$ $0.000$ $0.000$ $\bm{0.693}$ $0.237$ $0.056$ $0.012$ $0.002$ $0.000$ $0.000$
+$(0.000)$ $(0.000)$ $(0.053)$ $(0.030)$ $(0.016)$ $(0.005)$ $(0.001)$ $(0.000)$ $(0.000)$